### v2

# ScholaWrite: A Dataset of End-to-End Scholarly Writing Process

Abstract
Writing is a cognitively demanding task involving continuous decision-making, heavy use of working memory, and frequent switching between multiple activities. Scholarly writing is particularly complex as it requires authors to coordinate many pieces of multiform knowledge. To fully understand writers’ cognitive thought process, one should fully decode the end-to-end writing data (from individual ideas to final manuscript) and understand their complex cognitive mechanisms in scholarly writing. We introduce ScholaWrite dataset, the first-of-its-kind keystroke logs of an end-to-end scholarly writing process for complete manuscripts, with thorough annotations of cognitive writing intentions behind each keystroke. Our dataset includes LaTeX-based keystroke data from five preprints with nearly 62K total text changes and annotations across 4 months of paper writing. ScholaWrite shows promising usability and applications (e.g., iterative self-writing) for the future development of AI writing assistants for academic research, which necessitate complex methods beyond LLM prompting. Our experiments clearly demonstrated the importance of collection of end-to-end writing data, rather than the final manuscript, for the development of future writing assistants to support the cognitive thinking process of scientists. Our de-identified dataset, demo, and code repository are available on our project page111https://minnesotanlp.github.io/scholawrite/.
ScholaWrite: A Dataset of End-to-End Scholarly Writing Process
Linghe Wang††thanks: Equal contribution. Minhwa Lee Ross Volkov Luan Tuyen Chau Dongyeop Kang University of Minnesota {wang9257,lee03533,volko032,chau0139,dongyeop}@umn.edu
1 Introduction
Scholarly writing requires researchers to produce texts that concisely and precisely present novel findings and to follow a systematized structure and style by their targeted venue Jourdan et al. (2023); Kallestinova (2011); Bourekkache (2022). To address this need, researchers have leveraged large language models (LLMs) to develop intelligent support systems in several writing tasks, such as revision process Du et al. (2022b); Kim et al. (2022) or feedback generation Liang et al. (2024); Lu et al. (2024a). However, a critical question arises: Are LLMs able to truly understand and generate writing as human scientists do?
LLMs are generally trained to progress autoregressively, generating text from left to right. In contrast, human writing typically involves multiple iterations of complex and non-linear cognitive actions to refine the main message Flower and Hayes (1981); Koo et al. (2023), as illustrated in Figure 1. This fundamental cognitive gap between human writing and language models necessitates a deeper understanding of human writing processes and the development of cognitively-aligned writing assistance tools, moving beyond basic auto-completion.
Although much literature has examined the cognitive processes used in general writing tasks Lindgren and Sullivan (2003); Hayes and Berninger (2014); Zhu et al. (2023); Wengelin and Johansson (2023), the specific cognitive patterns in scholarly writing remain understudied. Our work observes the end-to-end writing process of scholarly writing, inspired by the method of keystroke collections that have been a major methodology to observe individual writing processes Baaijen et al. (2012); Latif (2008); Leijten and Van Waes (2013) in cognitive science or a few recent works in NLP at a small scale Koo et al. (2023); Velentzas et al. (2024).
We believe our work (ScholaWrite) is the first attempt to present a keystroke corpus of scholarly writing with annotations of cognitive writing processes, which were collected over multiple months and produced by early-career researchers. Furthermore, we developed novel data collection and annotation systems using Chrome extension, which improves data accessibility to the public and enables real-time collection of LaTeX-based keystrokes from multiple researchers. We also present a comprehensive taxonomy of cognitive writing processes specific to the scholarly writing domain. Throughout several experiments with LLMs, ScholaWrite shows promising usability as a resource for enabling LLMs to understand the human cognitive process of scholarly writing and provide writing suggestions that are aware of writers’ cognitive behaviors.
Contributions
We present ScholaWrite, a curated dataset of nearly 62K LaTeX-based keystrokes that were turned into publications in the computer science domain, annotated by experts in linguistics and computer science. We develop a taxonomy of scholarly writing intentions, providing an overall understanding of how scholars tend to develop writing manuscripts. Experiment results showed that the Llama-8B model fine-tuned on the ScholaWrite dataset achieved the high linguistic quality of the final writing.
2 Related Work
2.1 Cognitive Theory of Writing Process
Research on human writing significantly shifted from focusing solely on the final written product to exploring the cognitive stages of various writing phases Diederich (1974); Perl (1979); Krapels (1990); MacArthur and Graham (2016).
Building on this process-oriented approach, Flower and Hayes (1981)’s cognitive writing theory outlines three key cognitive sub-processes grounded in empirical evidence: (1) planning; (2) translating; and (3) reviewing. These processes are complex, interconnected, and non-linear, with writers moving fluidly between them. Our work expands this model to develop our own taxonomy of more granular cognitive patterns of human writing, specifically in scholarly communication.
Koo et al. (2023) presented a taxonomy of the scholarly writing process based on keystroke collections of just a short research plan over a thirty-minute session. Inspired by this work, we extend their approach to a larger-scale collection of keystrokes over months, whose final written products turned into research publications. Reviewed by a linguistics expert, our comprehensive taxonomy is built on top of prior literature and encompasses end-to-end writing trajectories of the cognitive process in a scholarly domain.
2.2 Keystroke Loggers for Scholarly Writing
Keystroke logging tools such as Inputlog allow researchers to observe digital writing without interfering with the writer’s process Chan (2017); Johansson et al. (2010); Leijten and Van Waes (2013); Lindgren and Sullivan (2019). However, they tend to operate in a closed ecosystem, i.e., only running in MS Word. This makes the current tools less accessible, especially for the scientific communities who use other typesetting such as LaTeX. Moreover, it is not well-suited for collecting data from extended writing sessions.
Currently, such keystroke logging studies have rarely examined scholarly communication settings that use LaTeX, especially for scientific writing purposes. To address this gap, our work implemented novel several interconnected systems that securely collect keystrokes in real-time over months from Overleaf, a widely-used online LaTeX editor, store them without privacy issues, and retrieve them for annotating writing intentions at span level. Our system workflow provides a more natural environment for observing cognitive writing activities by multiple researchers over long periods and across multiple sections without interrupt their writing.
2.3 Datasets for Scholarly Writing
Previous work primarily focused on constructing datasets for understanding the writing process of scholarly domains, which vary in their content and scale. Much of the publicly available datasets analyzed linguistic style changes or grammatical edits during the revision process Du et al. (2022b); Jiang et al. (2022); Ito et al. (2019); Mita et al. (2022). Also, the datasets from D’Arcy et al. (2024); Jourdan et al. (2024); Kuznetsov et al. (2022) collected edits conditioned on feedback and peer review process, while Kobayashi et al. (2022); Narimatsu et al. (2021) constructed datasets for citation generations. Existing work is also limited to only specific sections of scientific papers (e.g., abstract, introduction) Du et al. (2022b); Mita et al. (2022). In comparison, we construct a dataset that covers all cognitive phases of the writing process from its beginning to the final written product, through long periods of the data collection process.
Furthermore, those existing corpora for scientific writing research are created from the comparison between multiple versions of full-length, final products retrieved from preprint databases Jiang et al. (2022); Du et al. (2022b); D’Arcy et al. (2024); Jourdan et al. (2024). As pointed out by Jourdan et al. (2023), the comparison of finalized outputs does not present how those manuscripts evolved. Our work addresses this problem by building a corpus of keystrokes each of which represents the progress of the publications (See Figure 2).
3 Dataset Construction
To preserve writers’ natural thought flow and collect authentic writing data, we developed a Chrome extension that records keystroke data in the background. After the study, an in-depth annotation process is conducted to interpret the intent behind each keystroke. For open research, we make all the code of our data collection and annotation framework publicly available: https://minnesotanlp.github.io/scholawrite/. We describe the detailed process of our data collection, annotation system design and taxonomy creation in the following sections.
3.1 Participant Recruitment
We recruited 10 graduate students222All of them currently attend a four-year university in the U.S. and are proficient in English. from the computer science department who were actively preparing academic manuscripts in Overleaf LaTeX editor333https://overleaf.com/ for submission to peer-reviewed conferences in AI and NLP. The data is collected from November 2023 to February 2024, totaling up to 4 months. Our study has been approved by the IRB institution of the authors. Please refer to Appendix A.1 for more detailed descriptions.
3.2 Chrome Extension for Keystroke Collection in Overleaf
We designed and implemented a Chrome extension (Figure 4(a)(a)), which enables the real-time collection of keystroke trajectories in the Overleaf platform. Participants can create their account credentials, and after logging into the system, the extension monitors the user’s keystrokes in the background silently, without disrupting the typical writing process.
When a key-up event fires in a browser, the extension collects the writer’s viewable texts in the code editor panel. When each of these actions444Example actions are (1) inserting a space/newline; (2) copy/paste; (3) undo/redo; (4) switching files and (5) scrolling a page. occurs, the extension uses ‘diff_match_patch’ package555https://github.com/google/diff-match-patch to generate an array of differences between two subsequent texts (i.e., Figure 3). Then, the extension will send the difference array along with other metadata (e.g., time stamp, author ID, file name, etc.) to the backend server. To prevent any issue of private data collection, the backend fetches only the IDs of the Overleaf projects that participants consented to share during the recruitment process and filters out participants’ keystroke data from any unauthorized projects666We used Google Sheet API to retrieve ID information that we collected during the recruitment process..
For any Overleaf project that consists of multiple LaTex files, which is common in collaborative scholarly writing, we also collected all keystrokes from subfiles associated with the main LaTeX file. Our comprehensive data collection process captures the end-to-end writing processes of the participating authors across all parts of Overleaf projects. This approach ensures that our dataset reflects the full scope of scholarly writing including edits made in auxiliary files such as files of each section, appendix, bibliography, etc.
All keystroke data stored in the database were processed for annotation and model training purposes. Here, raw keystrokes from the Chrome extension system are stored as arrays that represent the differences between subsequent keystrokes, along with accompanying metadata.
| 1st | Intention | Definition | An example action | Prop. |
| Planning | Idea Generation | Formulate and record initial thoughts and concepts. | writing down keywords of a paragraph beforehand (e.g., “..%[Comment out] main point: artifacts lack in human subjectivity..”) | 7.0% |
| Idea Organization | Select the most useful materials and demarcate those generated ideas in a visually formatted way. | Linking the generated ideas into a logical sequence and spacing out between ideas (e.g., “..% (1) need diff. stress testing…%%[Spacing] (2) exp. setup? ”) | 0.5% | |
[…]
| Implementation | Text Production | Translate their ideas into full languages, either from the writers’ language or borrowed sentences from an external source. | Generating subsequent sentences with the author’s own idea (e.g., “… GPT-4 (OpenAI, 2023) explains the data … Our approach is built on top of GPT-4…”) | 57.4% |
[…]
| Coherence | Logically link (1) any of the two or multiple sentences within the same paragraph; (2) any two subsequent paragraphs; or (3) objects to be consistent as a whole. | “Each comment was annotated by three different annotators , which and we achieved high inter-annotator agreement.” | 3.3% | |
| Clarity | Improve the semantic relationships between texts to be more straightforward and concise. | “..relevant studies have examined one of the several textual styles one aspect of texts, the formality, ….” | 11.5% | |
| Structural | Improve the flow of information by modifying the location of texts and objects. | “We calculate Pearson’s correlation for human alignment to compare the alignment between lexicon-based preferences and humans’ original preferences. First, we calculated the score from each human participant to compare the alignment between lexicon-based preferences and humans’ original preferences.” | 3.7% | |
| Linguistic Style | Modify texts with the writer’s writing preferences regarding styles and word choices, etc. | “We believe posit that …” | 1.6% | |
| Scientific Accuracy | Update or correct scientific evidence (e.g., numbers, equations) for more accurate claims. | “..Pearson’s correlation (0.780.68; p < 0.01)” | 0.7% | |
[…]
3.3 Writing Intention Annotation
To capture comprehensive writing behaviors, we designed a data collection process that involves running a Chrome extension in the background. However, keystroke data alone may be insufficient to fully understand the nuanced intentions behind each writing action. To bridge this gap, we implement an additional annotation step to decode the writer’s intent for each keystroke using a specialized annotation interface. This process is crucial for developing a detailed taxonomy of the end-to-end scholarly writing process, which serves as the basis for annotating the collected keystrokes. Further details about the annotation process can be found in Appendix B.
To perform annotations, we first processed those raw keystroke entries by file name, type of writing actions (e.g., add, delete), an array of differences between two subsequent texts, and line number in an Overleaf editor.
Annotation Interface
We developed a novel interface for the keystroke annotation process. The interface offers various modes for visualizing keystroke collections within each Overleaf project: by time, LaTeX file, and author (Figure 4(a)(b)).
In the right panel, annotators can interactively right-click to navigate the timeline of collected keystrokes in different colors (e.g., red for deletions and green for additions). They then identify a range of keystrokes whose underlying purpose or goal is similar when making edits during the writing process (e.g., generating outlines or correcting grammatical errors). They can manually enter the beginning and ending indices and annotate the span by selecting a pre-defined label of the relevant intention from the dropdown menu in the left panel. Furthermore, the interface shows the metadata information of each keystroke, along with the selected label.
Intention Taxonomy Construction
Our taxonomy (Table 1) builds on prior literature Flower and Hayes (1981); Du et al. (2022b); Koo et al. (2023). During annotations, we have identified 15 types of scholarly writing intentions from the five projects and further grouped them into three high-level categories according to Flower and Hayes (1981) - Planning, Implementation, and Revision.
Following Pustejovsky et al. (2017)’s iterative annotate-and-refine framework, we performed an iterative open coding for the collected keystrokes from the two Overleaf projects. Using the annotation interface, each annotator reviewed the first few writing keystrokes and identified the high-level processes based on (Flower and Hayes, 1981)’s initial taxonomy of cognitive writing (planning, implementation, and revision). Next, the annotators (1) determine the span of keystrokes that form a meaningful unit together (e.g., a clause or sentence) within the high-level process and (2) decide the underlying intention of edits. The annotators coded these spans independently, resolving conflicts through multiple in-depth discussions. Finally, a cognitive linguist participated in all discussions, refining the definition and scope of each writing intention.
To ensure the validity of our taxonomy, we followed the two standard practices Nickerson et al. (2013); Kundisch et al. (2021): (1) All elements should be mutually exclusive and (2) the taxonomy should be collectively comprehensive. To achieve this, we flagged any ambiguous keystroke spans that could be interpreted as representing multiple intentions. Through iterative discussion, annotators refined these spans’ characteristics and improved each category’s decision boundaries. Once the codebook and taxonomy were finalized, we updated intention annotations777The list of those intentions is then updated in the dropdown bar of the annotation interface. for the selected two projects and applied the taxonomy to the remaining projects. The final inter-annotator agreement is using a weighted F1 score in a multi-label, multi-class setting for a randomly selected window of 1K keystrokes. Please refer to Appendix B for more details about the entire annotation procedures.
4 ScholaWrite: A Dataset of Cognitive Process of Scholarly Writing
| Project | 1 | 2 | 3 | 4 | 5 |
| # Authors | 1 (3) | 1 (4) | 1 (3) | 1 (4) | 9 (18) |
| # keystrokes | 14,217 | 5,059 | 6,641 | 8,348 | 27,239 |
| # Words added | 17,387 | 23,835 | 7,779 | 12,448 | 57,511 |
| # Words deleted | 11,739 | 15,158 | 2,308 | 7,621 | 25,853 |
| 1 | 2 | 3 | 4 | 5 | |
| Idea Generation | 515 | 130 | 116 | 309 | 3255 |
| Idea Organization | 0 | 45 | 25 | 9 | 231 |
| Section Planning | 182 | 57 | 111 | 201 | 773 |
| Text Production | 9267 | 2438 | 5109 | 4478 | 14031 |
| Object Insertion | 583 | 383 | 62 | 486 | 1300 |
| Cross-reference | 141 | 112 | 13 | 292 | 458 |
| Citation Integration | 75 | 151 | 69 | 127 | 245 |
| Macro Insertion | 16 | 7 | 51 | 29 | 33 |
| Linguistic Style | 233 | 75 | 42 | 201 | 411 |
| Coherence | 422 | 242 | 126 | 193 | 1021 |
| Clarity | 1249 | 645 | 721 | 1180 | 3301 |
| Scientific Accuracy | 307 | 15 | 2 | 24 | 95 |
| Structural | 359 | 506 | 105 | 257 | 1042 |
| Fluency | 116 | 90 | 46 | 135 | 476 |
| Visual Formatting | 752 | 163 | 43 | 427 | 567 |
Summary
Our post-processed dataset, ScholaWrite, contains end-to-end writing trajectories with intentions annotated, for five Overleaf projects whose final product turned into arXiv preprints. The dataset consists of 61,504 arrays of keystroke changes with fine-grained annotations of writing intentions. We publicly release our post-processed data in Huggingface data cards.888https://huggingface.co/datasets/minnesotanlp/scholawrite Table 3 shows the overall statistics of the dataset.
Writing Intention Distributions
The writing intention distributions is illustrated in Figure 5(a). During the planning stage, idea generation is the most frequent intention. Text production dominates the implementation stage, and clarity becomes the primary focus during the revision stage. Since we defined the boundary between implementation and revision as the point where meaning changes occur, the high frequency of text production suggests that authors often continue revising their texts by introducing new topics into the documents. Table 3 shows the label distribution per project.
Flows Between Writing Intentions
We analyze the flow of writing intentions most likely to follow each other throughout the writing process, as presented in Table 4. Many intentions proceed to text production, where the idea generation has the highest probability of feeding into text production. Also, the reflexive relationship between text production and clarity suggests a constant loop of producing new texts and refining them, ensuring the precision and reasonability of their claims.
| Label | Subsequent label | Probability | |
| Idea Generation | Text Production | 0.52 | |
| Idea Organization | Idea Generation | 0.34 | |
| Section Planning | Text Production | 0.33 | |
| Text Production | Clarity | 0.20 | |
| Object Insertion | Text Production | 0.32 | |
| Citation Integration | Text Production | 0.37 | |
| Cross-reference | Text Production | 0.36 | |
| Macro Insertion | Idea Generation | 0.29 | |
| Fluency | Text Production | 0.30 | |
| Coherence | Text Production | 0.34 | |
| Clarity | Text Production | 0.35 | |
| Structural | Text Production | 0.27 | |
| Linguistic Style | Text Production | 0.29 | |
| Scientific Accuracy | Text Production | 0.34 | |
| Visual Formatting | Text Production | 0.25 |
Time-series Distributions of Writing Intentions
We calculated the average Wasserstein distance between each intention distribution and uniform distribution to assess how evenly intentions occur throughout the writing process. As shown in Figure 5(b), text production had the smallest distance, while macro insertion had the largest. This suggests that text production is spread throughout the writing, whereas macro insertion is concentrated in specific phases. A similar pattern appears in Figure 5(c), where text production is evenly distributed over time, and macro insertion peaks in the early and final stages of the writing process. Figures 6(a) and 6(b) demonstrate a similar pattern that text production of the implementation stage spreads evenly over time. Types of clarity and structural revisions tend to occur after the medium steps of writing. Please refer to Figures 9 to 12 in the Appendix for each of the five Overleaf projects.
5 Applications for Writing Assistance with ScholaWrite
We believe ScholaWrite can serve as a valuable resource for training language models or enhancing the cognitive writing process in future writing assistants for scholarly writing. To evaluate its usability, we performed several experiments where existing LLMs were trained to mimic the complex, non-linear scientific writing processes of human scholars. Specifically, we aimed to showcase the capabilities of LLMs trained on ScholaWrite in two scenarios:
(1) Predicting an author’s next writing intention (Section 5.1)
: The task is crucial for writing assistants to accurately assess the writer’s current status in context and predict the correct writing intention. This enables them to offer cognitively-appropriate writing suggestions that align with the writer’s needs.
(2) Iteratively generating scholarly writing actions from scratch (called Iterative Self-Writing), mirroring the human writing process (Section 5.2)
: This task focuses on how well the model trained on our dataset can replicate the actual iterative writing and thinking process of scholars, and whether the generated text achieves higher quality compared to LLM-prompted writing.
As illustrated in Figure 7, the first task (represented by the inner box labeled "Prediction") takes the "before" text from a keystroke pair and context prompt, and predicts the writing intention to apply for the subsequent actions. Once the intention is predicted, the second task generates the “after” text based on the predicted intention and given prompt (represented by the outer box labeled “Generation”). This process repeats iteratively until no further changes are made or a maximum number of iterations, such as 100, is reached.
We randomly split our dataset into train (80%) and test (20%) sets. For each intention label in the test set, we randomly select up to 300 keystroke entries, due to budget constraints. Note that the same train and test sets are applied to all models across all experiments. See Appendix D for a full description of the model training process.
5.1 Predicting Next Writing Intention
Setup & Metrics
We use BERT Devlin et al. (2019), RoBERTa Liu et al. (2019), and Llama3.1-8B-Instruct Dubey et al. (2024) as baselines, fine-tuning each on the ScholaWrite training set. For comparison, we also run GPT-4o OpenAI (2024) on the test set. The fine-tuning prompt included all possible labels with definitions, task instructions, the “before” text chunk, and the corresponding human-annotated intention label, asking the model to predict the intention label based on the “before-text”. Differences in prompts were limited to only task instructions (see Appendix D.3 for prompt details). To evaluate model performance, we used a weighted f-1 score999Weighted F-1 was chosen to address skewed label distribution (as shown in Figure 5(a) and Table 1)., comparing predictions to gold intention labels on the test set.
| BERT | RoBERTa | LLama-8B | GPT-4o | |
| Base | 0.04 | 0.02 | 0.12 | 0.08 |
| + SW | 0.64 | 0.64 | 0.13 | - |
Results
Table 5 presents the weighted F1 scores for predicting writing intentions across baselines and fine-tuned models. Regardless of the intricate nature of the task itself101010Each model is asked to predict the next intention by only giving the ‘before’ text. Human annotators, however, label the data by looking through multiple consecutive ‘before and after’ edits to determine the current intention. In addition, the next intention chosen by the author does not necessarily mean that it is the only correct intention., all models finetuned on ScholaWrite show an improved performance compared to their baselines. BERT and RoBERTa achieved the most improvement, while LLama-8B-Instruct showed a modest improvement after fine-tuning. As detailed in Appendix D, more training epochs and encoder-decoder architectures of BERT variants are assumed to be the reason for significant improvement compared to LLMs. This aligns with findings from Grasso and Locci (2024); Yu et al. (2023), which shows that RoBERTa and BERT can often match or even outperform LLMs for the text classification tasks. Those results demonstrate the effectiveness of our ScholaWrite dataset to align language models with writers’ intentions.
5.2 Iterative Self-Writing
Setups
Iterative self-writing involves two subtasks (Figure 7): (1) next intention prediction and (2) "after-text" generation. For intention prediction, we use fine-tuned Llama3.1-8B-Instruct on the ScholaWrite training set (“Llama-8B-SW-pred”), compared to vanilla models (Llama-8B-Instruct, and GPT-4o) from Section 5.1 as baselines. Please refer to Figure 13 in Appendix D for the detailed pipeline.
For “after-text” generation, another vanilla Llama3.1-8B-Instruct is fine-tuned on the ScholaWrite training set (“Llama-8B-SW-gen”), with baselines including vanilla Llama3.1-8B-Instruct and GPT-4o. The fine-tuning prompt included task instructions, a verbalizer derived from human-annotated labels, and the “before” text. The model’s task is to generate “after-text” given the verbalizer and “before” text. To maintain consistency across outputs, prompts were adjusted for each model, with differences mainly in task instructions (see Appendix D.3 for the prompt templates).
Seed documents were derived from four award-winning NLP papers spanning different topics Zeng et al. (2024); Lu et al. (2024b); Du et al. (2022a); Etxaniz et al. (2024). Listing LABEL:table:seed-entry presents the LaTeX-formatted abstract of Du et al. (2022a) as an example seed text for initiating the iterative self-writing experiment.
6 Conclusion and Future Work
We introduce ScholaWrite, the first-of-its-kind dataset that captures the end-to-end scholarly writing process, consisting of nearly 62K LaTeX keystrokes collected via our custom-built Chrome extension. This dataset, sourced from ten graduate students with varying levels of scientific writing expertise, is further enriched with expert annotations based on a novel taxonomy of cognitive writing intentions inspired by Flower and Hayes (1981); Koo et al. (2023). Our study demonstrates the value of ScholaWrite as a resource for advancing the cognitive capabilities of large language models (LLMs) and developing cognitively-aligned writing assistants, enabling them to emulate the complex, non-linear, and intention-driven nature of human writing.
Future work includes expanding the dataset to diverse academic fields, authors, and collaborative projects, thus enabling models to generalize beyond fact recall to emulate human decision-making and reasoning in more realistic academic environments. Additionally, integrating advanced memory architectures and lifelong learning techniques could further enhance LLMs’ ability to adapt dynamically to evolving writing intentions and produce coherent, high-quality scholarly outputs.
Limitations and Ethical Considerations
We acknowledge several limitations in our study. First, the ScholaWrite dataset is currently limited to the computer science domain, as LaTeX is predominantly used in computer science journals and conferences. This domain-specific focus may restrict the dataset’s generalizability to other scientific disciplines. Future work could address this limitation by collecting keystroke data from a broader range of fields with diverse writing conventions and tools, such as the humanities or biological sciences. For example, students in humanities usually write book-length papers and integrate more sources, so it could affect cognitive complexities.
Second, our dataset includes contributions from only 10 participants, resulting in five final preprints on arXiv. This small-to-medium sample size is partly due to privacy concerns, as the dataset captures raw keystrokes that transparently reflect real-time human reasoning. To mitigate these concerns, we removed all personally identifiable information (PII) during post-processing and obtained full IRB approval for the study’s procedures. However, the highly transparent nature of keystroke data may still have discouraged broader participation. Future studies could explore more robust data collection protocols, such as advanced anonymization or de-identification techniques, to better address privacy concerns and enable larger-scale participation. We also call for community-wise collaboration and participation for our next version of our dataset, ScholaWrite 2.0 and encourage researchers to contact authors for future participation.
Furthermore, all participants were early-career researchers (e.g., PhD students) at an R1 university in the United States. Expanding the dataset to include senior researchers, such as post-doctoral fellows and professors, could offer valuable insights into how writing strategies and revision behaviors evolve with research experience and expertise. Despite these limitations, our study captured an end-to-end writing process for 10 unique authors, resulting in a diverse range of writing styles and revision patterns. The dataset contains approximately 62,000 keystrokes, offering fine-grained insights into the human writing process, including detailed editing and drafting actions over time. While the number of articles is limited, the granularity and volume of the data provide a rich resource for understanding writing behaviors. Prior research has shown that detailed keystroke logs, even from small datasets, can effectively model writing processes Leijten and Van Waes (2013); Guo et al. (2018); Vandermeulen et al. (2023). Unlike studies focused on final outputs, our dataset enables a process-oriented analysis, emphasizing the cognitive and behavioral patterns underlying scholarly writing.
Third, collaborative writing is underrepresented in our dataset, as only one Overleaf project involved multiple authors. This limits our ability to analyze co-authorship dynamics and collaborative writing practices, which are common in scientific writing. Future work should prioritize collecting multi-author projects to better capture these dynamics. Additionally, the dataset is exclusive to English-language writing, which restricts its applicability to multilingual or non-English writing contexts. Expanding to multilingual settings could reveal unique cognitive and linguistic insights into writing across languages.
Fourth, due to computational and cost constraints, we evaluated the usability of the ScholaWrite dataset with a limited number of LLMs and hyperparameter configurations. As shown in Table 5, the Llama-8B-Instruct model demonstrated only marginal improvements after fine-tuning on our dataset. This underscores the need for future research to explore advanced techniques, such as fine-grained prompt engineering, to better align LLM outputs with human writing processes. Specifically, optimizing prompts with clearer contextual guidance (e.g., "before-text" and intention label definitions) may significantly enhance model performance.
Finally, the human evaluation process in Section LABEL:sec:appendix:human-eval was determined as exempt from IRB review by the authors’ primary institution, while the data collection using our Chrome extension program was fully approved by the IRB at our institution. Importantly, no LLMs were used during any stage of the study, except for grammatical error correction in this manuscript.
Acknowledgements
This work was supported by the research gift from Grammarly. We thank Anna Martin-Boyle and Ryan Koo for their insights and the data collection tool based on their earlier version of this paper. We thank the participants of our data collection for their valuable time, effort, and contributions, which were essential to the success of this research. We also thank Minnesota NLP group members and anonymous reviewers for providing us with valuable feedback and comments on the paper draft.
Appendix A More About Data Collection Process
A.1 Participant Recruitment & Demographics
We recruited ten graduate students in the computer science department who actively prepared their manuscripts in Overleaf, an online LaTeX editor, and who aimed to submit their manuscripts to peer-reviewed conferences. We held a consent procedure with each participant through a 30-minute virtual meeting remotely. After the consent process, we installed on their computers a Chrome extension program that we designed and implemented only for this study and asked for the ID number of only the Overleaf projects that the participants agreed to share as their manuscripts. We provided all participants with a $100 Amazon gift card per project, which could be divided among the authors involved in the project. Our data collection process is approved by the IRB of the authors’ primary institution.
All ten participants of our study are graduate students who currently study computer science domain at an accredited university in the United States. Out of the ten, two of them identified themselves as native English speakers, and the remaining participants identified themselves as proficient in English in terms of writing. Also, two of the ten participants attained a Master of Science degree in Computer Science with several publication experiences, and the remaining eight of them are currently PhD students with extensive research experiences.
A.2 Technical Details of System Implementations
We explain the technical implementation details of the two systems for the data collection process. For the Chrome extension, we implemented a backend application using Flask and Python and stored all keystroke data in the MongoDB database.
For the annotation interface, we used HTML/CSS and JavaScript for the client side and Flask for the backend. All data for the annotation interface was retrieved from the MongoDB database used in the Chrome extension system.
A.3 Data Post-Processing
For use during the annotation phase, each keystroke entry from the raw collection includes the following fields: (1) a valid file name; (2) a valid writing action that triggered keystroke logging (e.g., copy, paste, typing, etc.); (3) a valid array of differences to enable visualization of writing trajectories; and (4) the line numbers in the Overleaf editor. Data entries annotated with a valid intention label (i.e., labels except ‘artifact’) and having a difference array length of fewer than or equal to are then used for model training.
Regarding the additional postprocessing for public use, we took the following steps to post-process our data with the annotations to promote the usability of our dataset and prevent any privacy issues. For the annotation data, we only include the keystroke changes, anonymized project ID, timeframe information, and anonymized author’s name (e.g., 0, 1, 2, etc.) from the metadata. Then, we extract the before and after texts from the differences array. We also include the annotated intention label for each entry.
[…]
Appendix B More About Writing Intention Annotations & Taxonomy
B.1 Annotator Recruitment
Due to privacy concerns, we did not hire external freelancers with expertise, rather the two corresponding authors of this paper annotated the data are graduate students who possess extensive scholarly writing experiences in natural language processing and data annotation skills. The raw keystroke data collected by our Chrome extension could potentially contain personally identifiable information, such as specific content edits or metadata that could reveal the identity of the authors. To ensure the confidentiality and ethical handling of sensitive information, we restricted access to the data to the authors only. This annotation process was also authorized by the IRB of the authors’ institution. Please note that the final dataset which will be released publicly is ensured not to contain any PII information through several sophisticated post-processing steps.
B.2 Detailed Annotation Process
The two authors (or annotators) collaborated with a cognitive linguist to develop a codebook and review the results of the annotations. Also, the two annotators conducted an iterative open coding approach to identify several unique writing intentions from keystrokes and developed a codebook of intention labels (“ground-truth labels”) within each high-level process (Planning, Implementation, and Revision) based on the findings from Flower and Hayes (1981); Koo et al. (2023). Using this codebook, those annotators re-labeled each span of keystrokes with the corresponding label during the annotation process.
The annotators were fully informed about all the labels and had complete access to them when annotating each data point. The annotation process for all the labels is the same: First, they view through multiple consecutive data points and identify which high-level label occurs (e.g., Planning, Implementation, or Revision). Once annotators have identified the current high-level label, attempting to identify where it ends. Then, they decide on the low-level label within the high-level label (e.g., idea generation or organization under the Planning stage, etc.). Finally, they identify the interval for low-level labels and annotate data points in the interval with the identified low-level label. If a keystroke does not deliver any insight, then label it as an ‘artifact.’
We calculated inter-annotator agreement using the weighted F1 score in a multi-label, multi-class setting, which is suitable for our complex annotation schema involving multiple labels per instance. The weighted F1 score achieved was 0.71, indicating a high level of agreement between the annotators.
Appendix C More About the Taxonomy of Scholarly Writing Process
During the planning stage, the writers engage in a process of generating and organizing raw ideas, arguments, or content structures that were not introduced in the previous trajectory. Based on the plan, the writers implement their plan by drafting full sentences and paragraphs and structuring the contents tangibly. At the same time, the writers enter the revision stage by improving the quality of their implemented sentences and LaTeX objects in terms of linguistic styles, format, or information accuracy. Particularly, spans of keystrokes whose intentions involved any changes but did not change the meaning of original texts are classified as Revision. For those edits that show changes in the meaning, we considered them as Implementation. Furthermore, if an author repeatedly adds, removes, and revises text back and forth until a sentence is completed, we consider this process as part of text production. Any subsequent changes made to the sentence after it is finished are considered revision.
Appendix D More about ScholaWrite Evaluation
D.1 Training environments
BERT & RoBERTa
We fine-tuned BERT and RoBERTa with the following hyperparameter setups: (1) a learning rate of ; (2) training batch size per device of 8; (3) evaluation batch size per device of 8; (4) the number of training epochs of 10; and (5) a weight decay of . For each model, it took approximately 3.5 hours on one NVIDIA RTX A6000.
Llama
For all experiments, we used baseline models of 4-bit quantized Llama-8B-Instruct111111https://huggingface.co/unsloth/Meta-Llama-3.1-8B-Instruct-bnb-4bit, using unsloth library121212https://github.com/unslothai/unsloth.
For the intention prediction task (Sec. 5.1), here are the hyperparameter setups for the Llama models: (1) only one epoch of training; (2) a weight decay of 0.01; (3) warm-up steps of 5; (4) learning rate of ; and (5) AdamW 8-bit optimizer. Due to computational constraints, we were able to run only one epoch for fine-tuning Llama models on our ScholaWrite dataset. For Llama-8B, it took approximately 8 hours on one RTX A5000.
For the ‘after’ text generation subtask from the iterative self-writing experiment (Sec. 5.2), we used the following hyperparameter setups for the fine-tuned Llama-8B-Instruct: (1) only one epoch of training; (2) a weight decay of 0.01; (3) warm-up steps of 10; (4) learning rate of ; and (5) AdamW 8-bit optimizer. Due to computational constraints, we were able to run only one epoch for fine-tuning Llama models on our ScholaWrite dataset. Also, it took approximately 12 hours on one NVIDIA L40s.
GPT-4o
We used the GPT-4o-2024-08-06 version for the classification and iterative writing experiments. The iterative writing with 100 iterations took approximately 1 hour on each seed.
D.2 Details About Finetuning Process
Next Intention Prediction (Sec 5.1)
The fine-tuning prompt consisted of task instructions, a list of all intention labels with their definitions, the “before” text, and the corresponding human-annotated intention label. To achieve optimal performance while minimizing memory usage, we employed QLoRA Dettmers et al. (2024) to fine-tune all linear modules of a 4-bit quantized Llama. During fine-tuning, we utilized the ‘train_on_response_only’ function provided by the unsloth library, which masks the task instructions, intention label definitions, and “before” text with -100s. This ensures the model is trained exclusively on the response portion of the fine-tuning prompt (i.e., the predicted intention label), without being influenced by the instructional components of the input. The model was fine-tuned for one epoch with a batch size of 2, 4 gradient accumulation steps, and the AdamW 8-bit optimizer.
Iterative Self-Writing (Sec 5.2)
The structure of the “after” text differs slightly to help the model learn where and what edits to make. We used the diff_match_patch library to generate a word-level difference array between the “before” and “after” texts. Special tokens (<same>, </same>, <del>, </del>, <add>, </add>) were added to the tokenizer, and the difference array was converted into text wrapped with these tokens. For example, given a “before” text of ‘‘Bad dog’’ and an “after” text of ‘‘Good dog’’, the difference array would be [(-1, ‘Bad’), (1, ‘Good ’), (0, ‘dog’)]. This is converted into: <del>Bad </del><add>Good </add><same>dog</same>. This transformation was applied only to the “after” text, while the “before” text remained as plain LaTeX text.
The fine-tuning prompt included task instructions, a verbalizer derived from human-annotated labels, and the “before” text. For fine-tuning, we used QLoRA Dettmers et al. (2024) to optimize all linear modules of a 4-bit quantized model while maintaining a small memory footprint. Additionally, the embed_tokens and lm_head modules were set as trainable and saved in the final checkpoint. To focus training on the response portion (the “after” text), we used the train_on_response_only function, masking the task instructions, verbalizer, and “before” text with -100s. This ensures the model learns to generate the "after" text without being influenced by instructional input. The model was trained for one epoch with a batch size of 1, 4 gradient accumulation steps, and the AdamW 8-bit optimizer.
During iterative writing, we performed 100 iterations, treating the model’s output under one intention as a single iteration. If the intention predicted by the classification model (fine-tuned Llama3.1-8B-Instruct, as described in Sec 5.1) matched the current predicted intention, the model was prompted to edit the text again. In this case, the newly generated output was not treated as final output in the iteration, and the iteration did not proceed. We moved to the next iteration only when the intention prediction model generated a different intention label than the previous one.
For model setups, we created three pairs of models for each prediction and generation subtasks as follows:
LLama-8B-SW: LLama-8B-Instruct fine-tuned on ScholaWrite dataset (Prediction) & LLama-8B-Instruct fine-tuned on ScholaWrite dataset (Generation), independently
LLama-8B-Instruct: Vanilla LLama-8B-Instruct (Prediction) & Vanilla LLama-8B-Instruct (Generation), independently
GPT-4o: GPT-4o inference (Prediction) & GPT-4o inference (Generation)
D.3 Fine-tuning Prompt Templates
D.3.1 Prediction Prompt for Llama-8B-Zero models
“Here are all the possible writing intention labels:
Idea Generation: Formulate and record initial thoughts and concepts.
Idea Organization: Select the most useful materials and demarcate those generated ideas in a visually formatted way.
Section Planning: Initially create sections and sub-level structures.
Text Production: Translate their ideas into full languages, either from the writers’ language or borrowed sentences from an external source.
Object Insertion: Insert visual claims of their arguments (e.g., figures, tables, equations, footnotes, itemized lists, etc.).
Cross-reference: Link different sections, figures, tables, or other elements within a document through referencing commands.
Citation Integration: Incorporate bibliographic references into a document and systematically link these references using citation commands.
Macro Insertion: Incorporate predefined commands or packages into a LaTeX document to alter its formatting.
Fluency: Fix grammatical or syntactic errors in the text or LaTeX commands.
Coherence: Logically link (1) any of the two or multiple sentences within the same paragraph; (2) any two subsequent paragraphs; or (3) objects to be consistent as a whole.
Structural: Improve the flow of information by modifying the location of texts and objects.
Clarity: Improve the semantic relationships between texts to be more straightforward and concise.
Linguistic Style: Modify texts with the writer’s writing preferences regarding styles and word choices, etc.
Scientific Accuracy: Update or correct scientific evidence (e.g., numbers, equations) for more accurate claims.
Visual Formatting: Modify the stylistic formatting of texts, objects, and citations.
Identify the most likely next writing intention of a graduate researcher when editing the following LaTex paper draft. Your output should only be a label from the list above.
## Input: before_text
## Output:
D.3.2 Generation Prompt for Llama-8B-Zero models
“You are a computer science researcher with extensive experience of scholarly writing. Here, you are writing a research paper in natural language processing using LaTeX languages.
You currently want to ‘‘put the verbalizer of the predicted intention label’’ (e.g., ‘‘initially create sections and sub-level structures’’ if the predicted label was section planning).
Below is the paper you have written so far. Please strictly follow the writing intention given above and insert, delete, or revise at appropriate places in the paper given below.
Your writing should relate to the paper given below. Do not generate text other than paper content. Do not describe the changes you are making or your reasoning.
## Input: before_text ”
D.3.3 Prediction Prompt for Llama-8B-SW models
“Here are all the possible writing intention labels:
Idea Generation: Formulate and record initial thoughts and concepts.
Idea Organization: Select the most useful materials and demarcate those generated ideas in a visually formatted way.
Section Planning: Initially create sections and sub-level structures.
Text Production: Translate their ideas into full languages, either from the writers’ language or borrowed sentences from an external source.
Object Insertion: Insert visual claims of their arguments (e.g., figures, tables, equations, footnotes, itemized lists, etc.).
Cross-reference: Link different sections, figures, tables, or other elements within a document through referencing commands.
Citation Integration: Incorporate bibliographic references into a document and systematically link these references using citation commands.
Macro Insertion: Incorporate predefined commands or packages into a LaTeX document to alter its formatting.
Fluency: Fix grammatical or syntactic errors in the text or LaTeX commands.
Coherence: Logically link (1) any of the two or multiple sentences within the same paragraph; (2) any two subsequent paragraphs; or (3) objects to be consistent as a whole.
Structural: Improve the flow of information by modifying the location of texts and objects.
Clarity: Improve the semantic relationships between texts to be more straightforward and concise.
Linguistic Style: Modify texts with the writer’s writing preferences regarding styles and word choices, etc.
Scientific Accuracy: Update or correct scientific evidence (e.g., numbers, equations) for more accurate claims.
Visual Formatting: Modify the stylistic formatting of texts, objects, and citations.
Identify the most likely next writing intention of a graduate researcher when writing the following LaTex paper draft. Your output should only be a label from the list above.
## Input: before_text
## Output:
D.3.4 Generation Prompt for Llama-8B-SW models
“You are a computer science researcher with extensive experience in scholarly writing. Here, you are writing a research paper in natural language processing using LaTeX.
You currently want to ‘‘put the verbalizer of the predicted intention label’’ (e.g., ‘‘initially create sections and sub-level structures’’ if the predicted label was section planning).
Below is the paper you have written so far. Given the paper information below and the corresponding scholarly writing intention, please revise or add to the text to fulfill this writing intention.
You may insert, delete, or revise text at appropriate places in the given paper.
Please provide a complete output. Do not generate text that is nonsensical or unrelated to the given paper information.
## Input: before_text
## Output:
D.3.5 Classification Prompt for GPT-4o
“You are a classifier that identify the most likely next writing intention. You will be given a list of all possible writing intention labels with definitions, and an in-progress LaTex paper draft written by a graduate student. Please strictly follow user’s instruction to identify the most likely next writing intention
Here is a list of all the possible writing intention labels with definitions:
Idea Generation: Formulate and record initial thoughts and concepts.
Idea Organization: Select the most useful materials and demarcate those generated ideas in a visually formatted way.
Section Planning: Initially create sections and sub-level structures.
Text Production: Translate their ideas into full languages, either from the writers’ language or borrowed sentences from an external source.
Object Insertion: Insert visual claims of their arguments (e.g., figures, tables, equations, footnotes, itemized lists, etc.).
Cross-reference: Link different sections, figures, tables, or other elements within a document through referencing commands.
Citation Integration: Incorporate bibliographic references into a document and systematically link these references using citation commands.
Macro Insertion: Incorporate predefined commands or packages into a LaTeX document to alter its formatting.
Fluency: Fix grammatical or syntactic errors in the text or LaTeX commands.
Coherence: Logically link (1) any of the two or multiple sentences within the same paragraph; (2) any two subsequent paragraphs; or (3) objects to be consistent as a whole.
Structural: Improve the flow of information by modifying the location of texts and objects.
Clarity: Improve the semantic relationships between texts to be more straightforward and concise.
Linguistic Style: Modify texts with the writer’s writing preferences regarding styles and word choices, etc.
Scientific Accuracy: Update or correct scientific evidence (e.g., numbers, equations) for more accurate claims.
Visual Formatting: Modify the stylistic formatting of texts, objects, and citations.
Identify the most likely next writing intention of a graduate researcher when writing the following LaTex paper draft. Your output should only be a label from the list above.
Here is LaTeX paper draft:
## Input: before_text
## Output:
D.3.6 Generation Prompt for GPT-4o
“ You are a computer science researcher with extensive experience of scholarly writing. Here, you are writing a research paper in natural language processing using LaTeX languages.
Your writing intention is to ‘‘put the verbalizer of the predicted intention label’’ (e.g., ‘‘initially create sections and sub-level structures’’ if the predicted label was section planning).
Below is the paper you have written so far. Please strictly follow the writing intention given above and insert, delete, or revise at appropriate places in the paper given below.
Your writing should relate to the paper given below. Do not generate text other than paper content. Do not describe the changes you are making or your reasoning. Do not include sidenotes. Your output should only be the paper draft in latex, without the “‘latex delimiters.
## Input: before_text
## Output:
D.4 Seed Documents for Iterative Self-Writing
We present the remaining three seed documents that we used for the iterative self-writing experiments, as shown in Listings LABEL:table:seed-entry-johnny to LABEL:table:seed-entry-latxa.