### v4

# ScholaWrite: A Dataset of End-to-End Scholarly Writing Process

Abstract
Writing is a cognitively demanding activity that requires constant decision-making, heavy reliance on working memory, and frequent shifts between tasks of different goals. To build writing assistants that truly align with writers’ cognition, we must capture and decode the complete thought process behind how writers transform ideas into final texts. We present ScholaWrite, the first dataset of end-to-end scholarly writing, tracing the multi-month journey from initial drafts to final manuscripts. We contribute three key advances: (1) a Chrome extension that unobtrusively records keystrokes on Overleaf, enabling the collection of realistic, in-situ writing data; (2) a novel corpus of full scholarly manuscripts, enriched with fine-grained annotations of cognitive writing intentions. The dataset includes LaTeX-based edits from five computer science preprints, capturing nearly 62K text changes over four months; and (3) analyses and insights into the micro-dynamics of scholarly writing, highlighting gaps between human writing processes and the current capabilities of large language models (LLMs) in providing meaningful assistance. ScholaWrite underscores the value of capturing end-to-end writing data to develop future writing assistants that support, not replace, the cognitive work of scientists.
ScholaWrite: A Dataset of End-to-End Scholarly Writing Process
Khanh Chi Le††thanks: Equal Contribution, Linghe Wang∗, Minhwa Lee∗, Ross Volkov, Luan Tuyen Chau, Dongyeop Kang University of Minnesota {le000422, wang9257, lee03533, volko032, chau0139, dongyeop}@umn.edu
1 Introduction
Scientific writing is one of the most cognitively demanding forms of human communication. Researchers must articulate novel ideas with precision and structure, iteratively refining arguments, evidence, and phrasing over time (Jourdan et al., 2023; Kallestinova, 2011; Bourekkache, 2022). Unlike linear text generation, scholarly writing is a complex cognitive process, writers continually plan, draft, and revise across multiple stages and intentions (see Fig. 1). Yet, despite the growing integration of large language models (LLMs) into research workflows, our understanding of how human scientists actually write remains limited. Without such understanding, it is difficult to build writing assistants that genuinely align with the cognitive processes of their users.
Recent efforts in writing assistance have leveraged LLMs for tasks such as revision or feedback generation (Du et al., 2022b; Liang et al., 2024). However, these models largely operate autoregressively, producing text from left to right, while human writing unfolds through recursive, non-linear cycles of ideation, organization, and refinement (Flower and Hayes, 1981). This fundamental cognitive gap raises a key question: how can we empirically capture and model the end-to-end scholarly writing process so that future LLMs can support, rather than substitute, human cognition?
Existing studies have examined writing processes through revision logs or small-scale keystroke analyses (Leijten and Van Waes, 2013; Koo et al., 2023), but most focus on short writing tasks or version-level comparisons of finalized drafts (Du et al., 2022b; Jiang et al., 2022; D’Arcy et al., 2024). Such approaches miss the fine-grained temporal and cognitive dynamics that drive real-world scholarly writing over extended periods.
To bridge this gap, we present ScholaWrite, the first dataset capturing the end-to-end scholarly writing process, from initial drafts to final manuscripts, through in-situ keystroke logging and cognitive annotation. Our work contributes three main advances:
(1) A reusable tool and taxonomy for cognitive writing capture. We develop a Chrome extension that unobtrusively records keystrokes within Overleaf, the dominant platform for collaborative scientific writing. Paired with a taxonomy of fifteen cognitive writing intentions, this tool enables large-scale, domain-specific capture of authentic writing behaviors (§3).
(2) The ScholaWrite dataset. Using this system, we collect over 61K keystroke-based text changes from five Computer Science preprints written by graduate researchers over four months. Each keystroke is annotated with its underlying cognitive intention (e.g., idea generation, clarity improvement), producing a richly structured corpus for modeling writing cognition (§4).
(3) Empirical insights into human-LLM gaps. Our analyses reveal that scholarly writing is highly non-linear and multi-intentional: more than half of all writing sessions involve three or more intertwined intentions. Writers dedicate most time to producing and revising text for clarity and coherence, while cognitively heavy tasks, such as idea organization or visual structuring, occur in fewer but longer, more focused sessions. Evaluations of current LLMs (GPT-5, Qwen) show they can imitate surface-level edits but fail to predict next intentions or sustain cognitively complex revisions, underscoring the limits of current models as cognitively aligned writing partners (§5 and §6).
2 Related Work
Cognitive Theory of Writing Process
Research on writing has shifted from analyzing final texts to examining cognitive processes across writing phases (Diederich, 1974; Krapels, 1990; MacArthur and Graham, 2016). Building on this process-oriented approach, Flower and Hayes (1981) outline three core sub-processes: planning, translating, and reviewing. These dynamic, non-linear stages inform our work, where we extend this model into a finer-grained taxonomy of cognitive writing patterns in scholarly communication. Koo et al. (2023) proposed a taxonomy of scholarly writing based on keystroke data from short, 30-minute research-plan sessions. Expanding on this, we collect larger-scale, longitudinal keystroke data spanning months and culminating in published research. Expert-reviewed and grounded in prior theory, our taxonomy captures end-to-end cognitive trajectories of scholarly writing.
Keystroke Loggers for Scholarly Writing
Keystroke logging tools (e.g., Inputlog) enable unobtrusive observation of digital writing (Chan, 2017; Johansson et al., 2010; Leijten and Van Waes, 2013; Lindgren and Sullivan, 2019). Yet, most operate in closed ecosystems like MS Word and are ill-suited for LaTeX-based academic writing or extended sessions. To address this, we built systems that securely record real-time keystrokes from Overleaf—an online LaTeX editor—over months while preserving privacy. This workflow supports span-level annotation of writing intentions, providing a natural setting to study long-term cognitive writing activities across document sections.
Datasets for Scholarly Writing
Publicly available datasets from previous work tend to track linguistic style changes or grammatical edits during revision (Du et al., 2022b; Jiang et al., 2022; Ito et al., 2019; Mita et al., 2022), while others capture edits based on feedback and peer review (D’Arcy et al., 2024; Jourdan et al., 2024; Kuznetsov et al., 2022) or citation generation (Kobayashi et al., 2022; Narimatsu et al., 2021). Also, most focus on specific sections of papers (e.g., abstracts, introductions) (Du et al., 2022b; Mita et al., 2022). In contrast, our dataset covers all cognitive phases of writing over an extended period and all sections. Furthermore, those existing corpora often compare multiple versions of final manuscripts from preprint databases (Jiang et al., 2022; Du et al., 2022b; D’Arcy et al., 2024; Jourdan et al., 2024), missing how those manuscripts evolved (Jourdan et al., 2023). To address this, we build a keystroke-based corpus that captures real-time progression of publications.
3 Data Collection and Intent Annotation
To capture scholarly writing as it naturally unfolds, we developed a system that records the end-to-end writing process, from individual keystrokes to their cognitive interpretation, without interrupting authors’ workflow. This section introduces the design of our Chrome extension for Overleaf-based data collection, the participant recruitment protocol, and the annotation pipeline that underpins our taxonomy of cognitive writing intentions.
3.1 Keystroke Collection in Overleaf
We built a custom Chrome extension (Appendix Figure 8) that unobtrusively logs real-time keystroke trajectories from Overleaf. The extension operates in the background once participants consent and authenticate via unique credentials. Every time a key-up event occurs, the system captures the visible text within the user’s Overleaf editor, compares it with the previous version using Google’s diff_match_patch algorithm, and stores the resulting text differences alongside metadata such as timestamp, file name, and author ID. To protect privacy, only project IDs pre-approved through participant consent were collected; any non-consented Overleaf projects were automatically filtered out. All data were stored on a secure institutional server and anonymized before analysis. Appendix A.2 provides further implementation details.
3.2 Participant Recruitment
We recruited 10 graduate students in computer science at an R1 university in the United States, each actively writing manuscripts for peer-reviewe AI and NLP conferences using Overleaf. The collection period spanned four months (November 2023–February 2024), capturing writing sessions over the full lifecycle of manuscript preparation. All participants were proficient in English and provided informed consent under IRB-approved protocols (Appendix A.1). Each participant installed the Chrome extension on their personal device and continued their usual writing activities, allowing us to gather naturalistic data without external disruption.
3.3 Writing Intention Annotation
To uncover the cognitive mechanisms underlying scholarly writing, we annotate every collected keystroke with its corresponding writing intention. This annotation process transforms raw writing traces into interpretable cognitive data, forming the foundation of the ScholaWrite taxonomy and subsequent analyses. Each keystroke record contains metadata, such as file name, type of action, text differences between two states, and line number in the Overleaf editor, which allows precise reconstruction of how writers iteratively develop and refine their manuscripts. Please see Appendix B for further annotation details.
Annotation Interface
We developed an interactive annotation interface that visualizes writing activity over time and across authors and files (Figure 9 in Appendix). Annotators can navigate a timeline of keystrokes, examine spans of related edits (e.g., drafting a sentence or clarifying an argument), and assign one or more intention labels from a predefined taxonomy via a dropdown menu. This design allows annotators to interpret not only what changes were made, but why they occurred, linking textual behavior to underlying cognitive intent.
| 1st | Intention | Definition | An example action | Prop. |
| Planning | Idea Generation | Formulate and record initial thoughts and concepts. | writing down keywords of a paragraph beforehand (e.g., “..%[Comment out] main point: artifacts lack in human subjectivity..”) | 7.0% |
| Idea Organization | Select the most useful materials and demarcate those generated ideas in a visually formatted way. | Linking the generated ideas into a logical sequence and spacing out between ideas (e.g., “..% (1) need diff. stress testing…%%[Spacing] (2) exp. setup? ”) | 0.5% | |
[…]
| Implementation | Text Production | Translate their ideas into full languages, either from the writers’ language or borrowed sentences from an external source. | Generating subsequent sentences with the author’s own idea (e.g., “… GPT-4 (OpenAI, 2023) explains the data … Our approach is built on top of GPT-4…”) | 57.4% |
[…]
| Revision | Fluency | Fix grammatical or syntactic errors in the text or LaTeX commands. | “We desiginingned several experiment setups for the LLM evaluations as described..” | 1.4% |
| Coherence | Logically link (1) multiple sentences within the same paragraph; (2) any two subsequent paragraphs; or (3) objects to be consistent. | “Each comment was annotated by three different annotators , which and we achieved high inter-annotator agreement.” | 3.3% | |
| Clarity | Improve the semantic relationships between texts to be more straightforward and concise. | “..relevant studies have examined one of the several textual styles one aspect of texts, the formality, ….” | 11.5% | |
| Structural | Improve the flow of information by modifying the location of texts and objects. | “..human alignment to compare the alignment between lexicon-based preferences and humans’ original preferences. First, we calculated the score from each human participant to compare the alignment between..” | 3.7% | |
| Linguistic Style | Modify texts with the writer’s writing preferences regarding styles and word choices, etc. | “We believe posit that …” | 1.6% | |
| Scientific Accuracy | Update or correct scientific evidence (e.g., numbers, equations) for more accurate claims. | “..Pearson’s correlation (0.780.68; p < 0.01)” | 0.7% | |
[…]
Intention Taxonomy Construction
Building on cognitive theories of writing (Flower and Hayes, 1981) and recent empirical studies on scholarly revision (Du et al., 2022b; Koo et al., 2023), we developed a taxonomy of 15 distinct writing intentions grouped into three overarching categories (Table 1). Following Pustejovsky et al. (2017), two annotators conducted iterative open coding on initial subsets of keystrokes to inductively identify recurring cognitive patterns. Each span of edits was treated as a meaningful writing unit, such as composing a sentence or improving phrasing, and assigned an appropriate intention label. Disagreements were resolved through multiple rounds of discussion with a cognitive linguist, who refined label definitions and boundaries to ensure coherence.
To validate the taxonomy, we applied two principles from taxonomy design (Nickerson et al., 2013; Kundisch et al., 2021): mutual exclusivity (each edit corresponds to one primary intention) and collective comprehensiveness (taxonomy covers all observed behaviors). After iterative refinement, we achieved a weighted F1 inter-annotator agreement of on a 1K-keystroke subset, confirming strong reliability. See details in Appendix B.
Post-processing
After annotation, we post-processed the data to improve usability and ensure privacy. All personal identifiers were removed, and keystrokes labeled as non-informative artifacts (e.g., scrolling, idle typing) were filtered out. Remaining entries were segmented into discrete intention spans suitable for analysis and model training. (See details in Appendix A.3).
4 ScholaWrite
ScholaWrite represents the first large-scale publicly available dataset capturing the cognitive dynamics of end-to-end scholarly writing. It links every keystroke to the author’s underlying writing intention, allowing direct observation of how ideas evolve into finalized text. Each annotated entry follows a standardized schema connecting text changes, contextual metadata, and intention labels, as shown below, illustrating a Fluency edit, in which the writer corrects “expct” to “aspect.”
4.1 Overview
The final ScholaWrite dataset contains 61,504 keystroke-based writing actions collected across five Overleaf projects, each of which culminated in an arXiv preprint authored by graduate researchers. Every action is annotated with a cognitive writing intention drawn from the 15-category taxonomy described in Section 3.3. Table 2 summarizes key statistics across projects. In total, the dataset captures over 62K text edits, encompassing more than 118K words added or deleted across multi-month writing trajectories. Each project contains thousands of temporally ordered keystrokes, providing a detailed chronicle of the writing process at the sentence, paragraph, and document levels.
4.2 Analytical Units: Capturing Writing Flow
Because the dataset records precise timestamps in milliseconds, we can reconstruct the temporal structure of scholarly writing. To analyze this structure, we define two complementary units: intention sessions and sessions.
Intention session.
An intention session is a continuous period during which the writer focuses on a single cognitive goal, such as producing text, improving clarity, or inserting a figure, before switching to another intention. These sessions typically last from a few seconds to several minutes, capturing a concentrated burst of focused cognitive activity. We apply two constraints for meaningful segmentation: (1) inactivity longer than 10 minutes marks the end of a session; (2) sessions shorter than 30 seconds are discarded to exclude trivial edits. This definition allows us to quantify how long writers sustain specific cognitive modes before transitioning to new tasks.
| Project | 1 | 2 | 3 | 4 | 5 |
| # Authors | 1 (3) | 1 (4) | 1 (3) | 1 (4) | 9 (18) |
| # Keystrokes | 14,217 | 5,059 | 6,641 | 8,348 | 27,239 |
| # Words added | 17,387 | 23,835 | 7,779 | 12,448 | 57,511 |
| # Words deleted | 11,739 | 15,158 | 2,308 | 7,621 | 25,853 |
Session.
In contrast, a session aggregates all writing actions within a continuous working period, regardless of how many intentions are involved. It represents the natural rhythm of scholarly writing, how authors weave together planning, drafting, and revising while working on a particular section or idea. Sessions end after 10 minutes of inactivity and must last at least one minute.
5 Analyses
ScholaWrite offers an unprecedented window into the dynamic, cognitive process of scholarly writing. Our analysis pursues two complementary goals: (1) to uncover how scholarly writing intentions evolve, interact, and shift across time, and (2) to inform the design of cognitively aligned writing assistants capable of supporting real-world research workflows. We focus on three research questions:
5.1 Which Writing Intentions Dominate?
We first examine which of the fifteen annotated writing intentions account for the greatest share of time and cognitive effort.
| Intention | % Time | Intention | % Time | ||
| Idea Generation | 4.8 | Fluency | 3.3 | ||
| Idea Organization | 0.6 | Coherence | 6.9 | ||
| Section Planning | 2.6 | Clarity | 13.4 | ||
| Text Production | 40.3 | Structural | 4.0 | ||
| Object Insertion | 8.4 | Linguistic Style | 2.6 | ||
| Citation Integration | 2.4 | Scientific Accuracy | 1.2 | ||
| Cross-reference | 2.4 | Visual Formatting | 6.1 | ||
| Macro Insertion | 0.8 |
Time Distribution across Intentions
Table 3 summarizes the average share of time spent per intention across five projects. Text Production dominates with 40.3% of total time, followed by Revision intentions – mainly Clarity, Coherence, and Visual Formatting – at 26.4%. In contrast, Planning activities (e.g., Idea Generation, Organization) take less than 5%.
These results suggest that scholars devote most of their time to articulating and refining text rather than explicit planning. The considerable effort spent on revision and formatting underscores that much of scholarly writing involves clarifying and visually shaping ideas, an aspect often overlooked in computational models of writing.
Session Duration and Cognitive Effort
We analyze the average duration of each intention session – a continuous period focused on one cognitive goal. Table 4 shows that Idea Organization sessions last the longest (3.7 min), followed by Object Insertion (2.9 min) and Visual Formatting (2.2 min), all showing high variability due to complex tasks like reorganizing structure or managing visuals.
In contrast, Text Production, though dominant in total time, consists of short, frequent bursts (mean 2.2 min). This contrast reveals two cognitive modes: (1) short, frequent drafting episodes, and (2) long, focused planning or visual composition that require sustained attention.
| Intention | Mean | SD | Max | Count |
| Idea Generation | 1.87 | 1.94 | 11.40 | 76 |
| Idea Organization | 3.71 | 4.58 | 13.79 | 7 |
| Section Planning | 1.37 | 1.57 | 9.36 | 38 |
| Text Production | 2.20 | 2.53 | 28.14 | 724 |
| Object Insertion | 2.88 | 3.05 | 21.61 | 108 |
| Citation Integration | 1.70 | 1.53 | 7.53 | 34 |
| Cross-reference | 1.08 | 0.77 | 3.98 | 38 |
| Macro Insertion | 1.43 | 1.05 | 3.71 | 7 |
| Fluency | 0.66 | 0.18 | 1.06 | 10 |
| Coherence | 1.46 | 1.71 | 9.62 | 55 |
| Clarity | 1.64 | 1.61 | 9.53 | 243 |
| Structural | 1.50 | 1.43 | 8.42 | 53 |
| Linguistic Style | 2.05 | 2.55 | 10.81 | 24 |
| Scientific Accuracy | 1.97 | 1.60 | 6.44 | 20 |
| Visual Formatting | 2.22 | 2.66 | 14.25 | 87 |
5.2 How Are Writing Intentions Intertwined?
Scholarly writing rarely unfolds linearly. Instead, it involves recursive alternations between complementary intentions, writers generate ideas, structure them, produce text, and immediately refine it. As seen in Figure 1 (and other projects in Appendix Figure 14), the intertwined processes of writing reveals dense bidirectional flows across intentions.
Transition Patterns
The transition probability matrix (Figure 2) shows that more than half of all Planning actions (e.g., Idea Generation) are immediately followed by Text Production. Conversely, Idea Organization often leads back to Idea Generation, suggesting that structuring ideas can prompt new content creation. Within Implementation, most transitions also return to Text Production (>30%), except for Macro Insertion, which frequently precedes Idea Generation. In Revision, strong recursive loops emerge, especially among Clarity, Fluency, and Coherence, highlighting how writers iteratively refine language and argumentation.
Overall, the results indicate that planning, implementation, and revision are not distinct stages but interdependent cycles continually revisited throughout the writing process.
Multitasking and Alternation Patterns
Writing sessions are highly multitasking in nature. As shown in Figure 3, 57% of all sessions involve three or more distinct intentions, reflecting frequent switching between cognitive modes within a single sitting. This dynamic interplay reveals how writers manage multiple goals in parallel.
Frequent 6-gram intention sequences (Appendix Table 9) confirm this recursive pattern. The most common alternating pairs include: (Text Production, Clarity), (Text Production, Object Insertion), and (Idea Generation, Idea Organization). These short feedback loops illustrate the micro-structure of scientific writing: writers repeatedly produce, adjust, and reorganize – blending idea formation with textual execution.
| Intention | Early | Late | Trend |
| Idea Gen. | 1.6 (23) | 2.5 (11) | longer, fewer |
| Idea Org. | 2.6 (2) | — | early-only |
| Sect. Plan. | 1.6 (13) | 1.0 (17) | more, shorter |
| Text Prod. | 2.1 (221) | 2.0 (227) | stable |
| Obj. Insert. | 2.4 (28) | 3.4 (42) | up both |
| Cite Int. | 1.4 (16) | 1.8 (6) | fewer later |
| Cross-ref. | 0.7 (8) | 1.2 (18) | more later |
| Clarity | 1.5 (43) | 1.3 (99) | more, quicker |
| Structural | 2.4 (8) | 1.4 (29) | more later |
| Ling. Style | 3.5 (3) | 1.7 (12) | concise edits |
| Vis. Form. | 3.6 (16) | 1.7 (43) | quick fixes |
5.3 Are writing patterns stable throughout the process, or do they change across phases?
Scholarly writing is not static, it evolves as authors move from idea formation to refinement. Beyond the aggregate trends observed earlier, we examine how writing intentions shift across time by comparing the early phase (first third of project duration) with the late phase (final third).
Which intents dominate at different stages?
Figure 4(a) shows how writing time is allocated across phases. Planning intentions are more frequent early (9.1% vs. 5.4%), reflecting front-loaded structuring. Within Implementation, Citation Integration is higher early, while Cross-referencing increases late as authors add visuals and refine internal links. Revision grows sharply in the late phase (46.7% vs. 29.6%), consistent with end-stage polishing for coherence and presentation.
Table 5 summarizes session-level differences. In Planning, Idea Generation and Organization sessions decrease in count but lengthen, suggesting deeper cognitive engagement later. Section Planning rises as writers reorganize material for submission. For Revision session counts increase substantially, though durations shorten, indicating rapid, localized edits characteristic of final revisions.
How are intentions intertwined across stages?
Figure 15 in Appendix compares transition probabilities across phases. Early writing features strong flows into Implementation, whereas late writing shifts toward Revision, marking the transition from building content to refining it. Some intention paths also evolve — for example, transitions from Macro Insertion lead to Idea Generation early but to Section Planning late, as layout decisions become more prominent. Writers also juggle more intentions later in the process. As shown in Figure 4(b), 19.4% of late sessions include over five distinct intentions (vs. 14.8% early), reflecting increased multitasking and integration of planning, drafting, and revising.
| BERT | RoBERTa | Llama-8B | GPT-4o | |
| Base | 0.04 | 0.02 | 0.12 | 0.08 |
| + SW | 0.64 | 0.64 | 0.13 | - |
6 Can LLMs Support Human Writing?
Scientific writing is deeply cognitive, non-linear, and phase-dependent. Writers continuously shift between intents, and multitasking across intertwined intentions. For LLMs to genuinely assist this process, they must (i) infer what the writer is doing or about to do, and (ii) provide support that aligns with the writer’s actual cognitive intent. We evaluate these capabilities through two tests using ScholaWrite:
6.1 Predicting Next Writing Intention
Because writing involves frequent task-switching, an effective assistant should anticipate the writer’s next move. We test whether models can predict the upcoming writing intention based on the current text state.
Setup.
We use the “before-text” from each keystroke pair and prompt the model to predict the next writing intention. Models evaluated include BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019), Llama3.1-8B-Instruct (Dubey et al., 2024), and GPT-4o, with fine-tuned versions train on our dataset. Performance is measured using weighted F1 to account for label imbalance. See details in Appendix G.2.
Findings.
Base models perform poorly (F1 0.12), indicating that current LLMs cannot infer writing intent from text alone. Fine-tuning on ScholaWrite improves performance dramatically (to 0.64 for BERT/RoBERTa), showing that cognitively annotated writing data helps models anticipate human writing behavior, a key step toward adaptive, context-aware assistance. Training details and explanations for differences across models are reported in Appendix F.1.
6.2 Output Alignment across Writing Sessions
To assist writers effectively, a model must not only infer their intentions but also generate text aligned with those goals. We evaluate LLMs under two settings: (1) single-intention sessions, focused on one cognitive goal, and (2) multi-intention sessions, where multiple intentions intertwine.
Setup.
For each session, models receive the “before-text” and either a single intention label or the sequence of intentions observed in that session, then generate the corresponding “after-text.” We use GPT-5 (OpenAI, 2025) as the base model and evaluate alignment with human writing using BERTScore-F1 (Zhang et al., 2019). Results for other models (e.g., Qwen (Yang et al., 2024)) and metrics (e.g., Levenshtein distance) appear in the Appendix F.2 and show consistent patterns.
Findings.
Model alignment decreases as sessions grow longer or cognitively more complex (Figure 5). As shown in Figure 6, GPT-5 (and Qwen) perform best on Revision intentions (e.g., Clarity, Fluency) but struggle with Planning and Implementation, which demand deeper reasoning and structural organization. Alignment further drops when multiple intentions co-occur within a session, indicating models’ limited ability to integrate overlapping cognitive goals. Overall, current LLMs excel at surface refinement but remain weak at sustaining the intertwined reasoning processes that drive human scholarly writing.
7 Conclusion and Future Work
This work introduces ScholaWrite, the first dataset to capture the end-to-end cognitive process of scholarly writing. Unlike prior work that centers only on final manuscripts, ScholaWrite reveals how writing unfolds through iterative, multi-intent, and phase-dependent behaviors. Our analyses uncover three key insights. First, scholarly writing is highly non-linear: authors continually alternate between planning, implementing, and revising, rather than progressing in discrete stages. Second, writing intentions evolve across phases, with early stages emphasizing idea development and late stages dominated by refinement and multitasking. Third, current LLMs, though proficient at surface-level revision, struggle to anticipate or integrate complex, cognitively demanding intentions, revealing a critical gap between human writing cognition and model capabilities.
ScholaWrite contributes: (1) a fine-grained lens on the cognitive dynamics of scholarly writing; (2) a benchmark for evaluating LLMs’ alignment with human writing intentions; and (3) a training resource for developing models that adapt to evolving, multi-intent writing behavior. Our Chrome extension enables unobtrusive, in-situ data capture in Overleaf, demonstrating a scalable and ethical method for collecting realistic writing data.
Future work will extend beyond keystroke logs to include pre-writing ideation, collaboration, and multimodal composition (e.g., figure and data integration). By tracing the full scholarly writing lifecycle, we aim to deepen our understanding of human cognitive patterns and to design AI writing assistants that not only complete text, but truly complement how humans think and write.
Limitations and Ethical Considerations
We acknowledge several limitations in our study. First, the ScholaWrite dataset is currently limited to the computer science domain, as LaTeX is predominantly used in computer science journals and conferences. This domain-specific focus may restrict the dataset’s generalizability to other scientific disciplines. Future work could address this limitation by collecting keystroke data from a broader range of fields with diverse writing conventions and tools, such as the humanities or biological sciences. For example, students in humanities usually write book-length papers and integrate more sources, so it could affect cognitive complexities.
Second, our dataset includes contributions from only 10 participants, resulting in five final preprints on arXiv. This small-to-medium sample size is partly due to privacy concerns, as the dataset captures raw keystrokes that transparently reflect real-time human reasoning. To mitigate these concerns, we removed all personally identifiable information (PII) during post-processing and obtained full IRB approval for the study’s procedures. However, the highly transparent nature of keystroke data may still have discouraged broader participation. Future studies could explore more robust data collection protocols, such as advanced anonymization or de-identification techniques, to better address privacy concerns and enable larger-scale participation. We also call for community-wise collaboration and participation for our next version of our dataset, ScholaWrite 2.0 and encourage researchers to contact authors for future participation.
Furthermore, all participants were early-career researchers (e.g., PhD students) at an R1 university in the United States. Expanding the dataset to include senior researchers, such as post-doctoral fellows and professors, could offer valuable insights into how writing strategies and revision behaviors evolve with research experience and expertise. Despite these limitations, our study captured an end-to-end writing process for 10 unique authors, resulting in a diverse range of writing styles and revision patterns. The dataset contains approximately 62,000 keystrokes, offering fine-grained insights into the human writing process, including detailed editing and drafting actions over time. While the number of articles is limited, the granularity and volume of the data provide a rich resource for understanding writing behaviors. Prior research has shown that detailed keystroke logs, even from small datasets, can effectively model writing processes (Leijten and Van Waes, 2013; Guo et al., 2018; Vandermeulen et al., 2023). Unlike studies focused on final outputs, our dataset enables a process-oriented analysis, emphasizing the cognitive and behavioral patterns underlying scholarly writing.
Third, collaborative writing is underrepresented in our dataset, as only one Overleaf project involved multiple authors. This limits our ability to analyze co-authorship dynamics and collaborative writing practices, which are common in scientific writing. Future work should prioritize collecting multi-author projects to better capture these dynamics. Additionally, the dataset is exclusive to English-language writing, which restricts its applicability to multilingual or non-English writing contexts. Expanding to multilingual settings could reveal unique cognitive and linguistic insights into writing across languages.
Finally, the human evaluation process in Section LABEL:sec:appendix:human-eval was determined as exempt from IRB review by the authors’ primary institution, while the data collection using our Chrome extension program was fully approved by the IRB at our institution. Importantly, no LLMs were used during any stage of the study, except for grammatical error correction in this manuscript.
Appendix A More About Data Collection Process
A.1 Participant Recruitment & Demographics
We recruited ten graduate students in the computer science department who actively prepared their manuscripts in Overleaf, an online LaTeX editor, and who aimed to submit their manuscripts to peer-reviewed conferences. We held a consent procedure with each participant through a 30-minute virtual meeting remotely. After the consent process, we installed on their computers a Chrome extension program that we designed and implemented only for this study and asked for the ID number of only the Overleaf projects that the participants agreed to share as their manuscripts. We provided all participants with a $100 Amazon gift card per project, which could be divided among the authors involved in the project. Our data collection process is approved by the IRB of the authors’ primary institution.
All ten participants of our study are graduate students who currently study computer science domain at an accredited university in the United States. Out of the ten, two of them identified themselves as native English speakers, and the remaining participants identified themselves as proficient in English in terms of writing. Also, two of the ten participants attained a Master of Science degree in Computer Science with several publication experiences, and the remaining eight of them are currently PhD students with extensive research experiences.
A.2 Technical Details of System Implementations
When a key-up event fires in a browser, the extension collects the writer’s viewable texts in the code editor panel111To prevent privacy concerns, the extension filters out keystroke data from any unauthorized Overleaf projects. Please see Appendix A.2 for more details.. When each of these actions222Example actions are (1) inserting a space/newline; (2) copy/paste; (3) undo/redo; (4) switching files and (5) scrolling a page. occurs, the extension uses ‘diff_match_patch’ package333https://github.com/google/diff-match-patch to generate an array of differences between two subsequent texts (i.e., Figure 7). Then, the extension will send the array along with metadata (e.g., time stamp, author ID, etc.) to the backend server.
For any Overleaf project that consists of multiple LaTeX files, we also collected all keystrokes from subfiles associated with the main LaTeX file. Our comprehensive data collection process captures the end-to-end writing processes of the participating authors across all parts of Overleaf projects. This approach ensures that our dataset reflects the full scope of scholarly writing including edits made in auxiliary files such as files of each section, appendix, bibliography, etc.
We explain the technical implementation details of the two systems for the data collection process. For the Chrome extension, we implemented a backend application using Flask and Python and stored all keystroke data in the MongoDB database.
For the annotation interface, we used HTML/CSS and JavaScript for the client side and Flask for the backend. All data for the annotation interface was retrieved from the MongoDB database used in the Chrome extension system.
Privacy Concerns
To prevent any issue of private data collection, we designed the backend of our Chrome extension to fetch only the IDs of the Overleaf projects that participants consented to share during the recruitment process and filter out participants’ keystroke data from any unauthorized projects. We used Google Sheet API to retrieve ID information that we collected during the recruitment process.
A.3 Data Post-Processing
For use during the annotation phase, each keystroke entry from the raw collection includes the following fields: (1) a valid file name; (2) a valid writing action that triggered keystroke logging (e.g., copy, paste, typing, etc.); (3) a valid array of differences to enable visualization of writing trajectories; and (4) the line numbers in the Overleaf editor. Data entries annotated with a valid intention label (i.e., labels except ‘artifact’) and having a difference array length of fewer than or equal to are then used for model training.
Regarding the additional postprocessing for public use, we took the following steps to post-process our data with the annotations to promote the usability of our dataset and prevent any privacy issues. For the annotation data, we only include the keystroke changes, anonymized project ID, timeframe information, and anonymized author’s name (e.g., 0, 1, 2, etc.) from the metadata. Then, we extract the before and after texts from the differences array. We also include the annotated intention label for each entry.
[…]
Appendix B More About Writing Intention Annotations & Taxonomy
B.1 Annotator Recruitment
Due to privacy concerns, we did not hire external freelancers with expertise, rather the two corresponding authors of this paper annotated the data are graduate students who possess extensive scholarly writing experiences in natural language processing and data annotation skills. The raw keystroke data collected by our Chrome extension could potentially contain personally identifiable information, such as specific content edits or metadata that could reveal the identity of the authors. To ensure the confidentiality and ethical handling of sensitive information, we restricted access to the data to the authors only. This annotation process was also authorized by the IRB of the authors’ institution. Please note that the final dataset which will be released publicly is ensured not to contain any PII information through several sophisticated post-processing steps.
B.2 Detailed Annotation Process
The two authors (or annotators) collaborated with a cognitive linguist to develop a codebook and review the results of the annotations. Also, the two annotators conducted an iterative open coding approach to identify several unique writing intentions from keystrokes and developed a codebook of intention labels (“ground-truth labels”) within each high-level process (Planning, Implementation, and Revision) based on the findings from Flower and Hayes (1981); Koo et al. (2023). Using this codebook, those annotators re-labeled each span of keystrokes with the corresponding label during the annotation process.
The annotators were fully informed about all the labels and had complete access to them when annotating each data point. The annotation process for all the labels is the same: First, they view through multiple consecutive data points and identify which high-level label occurs (e.g., Planning, Implementation, or Revision). Once annotators have identified the current high-level label, attempting to identify where it ends. Then, they decide on the low-level label within the high-level label (e.g., idea generation or organization under the Planning stage, etc.). Finally, they identify the interval for low-level labels and annotate data points in the interval with the identified low-level label. If a keystroke does not deliver any insight, then label it as an ‘artifact.’
We calculated inter-annotator agreement using the weighted F1 score in a multi-label, multi-class setting, which is suitable for our complex annotation schema involving multiple labels per instance. The weighted F1 score achieved was 0.71, indicating a high level of agreement between the annotators.
Appendix C More About the Taxonomy of Scholarly Writing Process
During the planning stage, the writers engage in a process of generating and organizing raw ideas, arguments, or content structures that were not introduced in the previous trajectory. Based on the plan, the writers implement their plan by drafting full sentences and paragraphs and structuring the contents tangibly. At the same time, the writers enter the revision stage by improving the quality of their implemented sentences and LaTeX objects in terms of linguistic styles, format, or information accuracy. Particularly, spans of keystrokes whose intentions involved any changes but did not change the meaning of original texts are classified as Revision. For those edits that show changes in the meaning, we considered them as Implementation. Furthermore, if an author repeatedly adds, removes, and revises text back and forth until a sentence is completed, we consider this process as part of text production. Any subsequent changes made to the sentence after it is finished are considered revision. Table 1 presents the comprehensive, complete definitions of each intention of end-to-end scholarly writing process, identified from ScholaWrite dataset.
Appendix D ScholaWrite dataset statistics
Table 8 shows the distribution of intention labels per Overleaf project from ScholaWrite.
| Label | Subsequent label | Probability | |
| Idea Generation | Text Production | 0.52 | |
| Idea Organization | Idea Generation | 0.34 | |
| Section Planning | Text Production | 0.33 | |
| Text Production | Clarity | 0.20 | |
| Object Insertion | Text Production | 0.32 | |
| Citation Integration | Text Production | 0.37 | |
| Cross-reference | Text Production | 0.36 | |
| Macro Insertion | Idea Generation | 0.29 | |
| Fluency | Text Production | 0.30 | |
| Coherence | Text Production | 0.34 | |
| Clarity | Text Production | 0.35 | |
| Structural | Text Production | 0.27 | |
| Linguistic Style | Text Production | 0.29 | |
| Scientific Accuracy | Text Production | 0.34 | |
| Visual Formatting | Text Production | 0.25 |
| 1 | 2 | 3 | 4 | 5 | |
| Idea Generation | 515 | 130 | 116 | 309 | 3255 |
| Idea Organization | 0 | 45 | 25 | 9 | 231 |
| Section Planning | 182 | 57 | 111 | 201 | 773 |
| Text Production | 9267 | 2438 | 5109 | 4478 | 14031 |
| Object Insertion | 583 | 383 | 62 | 486 | 1300 |
| Cross-reference | 141 | 112 | 13 | 292 | 458 |
| Citation Integration | 75 | 151 | 69 | 127 | 245 |
| Macro Insertion | 16 | 7 | 51 | 29 | 33 |
| Linguistic Style | 233 | 75 | 42 | 201 | 411 |
| Coherence | 422 | 242 | 126 | 193 | 1021 |
| Clarity | 1249 | 645 | 721 | 1180 | 3301 |
| Scientific Accuracy | 307 | 15 | 2 | 24 | 95 |
| Structural | 359 | 506 | 105 | 257 | 1042 |
| Fluency | 116 | 90 | 46 | 135 | 476 |
| Visual Formatting | 752 | 163 | 43 | 427 | 567 |
Figures 10 to 13 show several characteristics of human writing process, analyzed from ScholaWrite dataset: (1) Figure 10 - the average Wasserstein distance between each intention distribution and uniform distribution; (2) Figure 11 - distribution of labels over time; (3) Figure 12 for high-level intention distribution over time; and (4) Figure 13 for intention-wise writing activity distribution over time.
Appendix E More about Analyses
Figure 15 shows transition probability between pair of intentions in early stage and later stage.
Table 9 shows top 6-gram writing intention sequences by session coverage (%)
| 6-gram Sequence | Percentage |
| Text Production Clarity Text Production Clarity Text Production Clarity | 3.22 |
| Clarity Text Production Clarity Text Production Clarity Text Production | 3.22 |
| Text Production Object Insertion Text Production Object Insertion Text Production Object Insertion | 2.34 |
| Object Insertion Text Production Object Insertion Text Production Object Insertion Text Production | 2.34 |
| Idea Generation Idea Organization Idea Generation Idea Organization Idea Generation Idea Organization | 2.05 |
Appendix F More about Experiments
F.1 Predicting Next Writing Intention
F.1.1 Training environments
BERT & RoBERTa
We fine-tuned BERT and RoBERTa with the following hyperparameter setups: (1) a learning rate of ; (2) training batch size per device of 8; (3) evaluation batch size per device of 8; (4) the number of training epochs of 10; and (5) a weight decay of . For each model, it took approximately 3.5 hours on one NVIDIA RTX A6000.
Llama
For all experiments, we used baseline models of 4-bit quantized Llama-8B-Instruct444https://huggingface.co/unsloth/Meta-Llama-3.1-8B-Instruct-bnb-4bit, using unsloth library555https://github.com/unslothai/unsloth.
For the intention prediction task (Sec. 6.1), here are the hyperparameter setups for the Llama models: (1) only one epoch of training; (2) a weight decay of 0.01; (3) warm-up steps of 5; (4) learning rate of ; and (5) AdamW 8-bit optimizer. Due to computational constraints, we were able to run only one epoch for fine-tuning Llama models on our ScholaWrite dataset. For Llama-8B, it took approximately 8 hours on one RTX A5000.
GPT-4o
We used the GPT-4o-2024-08-06 version.
F.1.2 Details About Finetuning Process
The fine-tuning prompt included all possible labels with definitions, task instructions, the “before-text” chunk, and the corresponding human-annotated intention label, asking the model to predict the intention label based on the “before-text”. Differences in prompts were limited to only task instructions (see §G.3 for prompt details).
To achieve optimal performance while minimizing memory usage, we employed QLoRA (Dettmers et al., 2024) to fine-tune all linear modules of a 4-bit quantized Llama. During fine-tuning, we utilized the ‘train_on_response_only’ function provided by the unsloth library, which masks the task instructions, intention label definitions, and “before” text with -100s. This ensures the model is trained exclusively on the response portion of the fine-tuning prompt (i.e., the predicted intention label), without being influenced by the instructional components of the input. The model was fine-tuned for one epoch with a batch size of 2, 4 gradient accumulation steps, and the AdamW 8-bit optimizer.
F.1.3 Prompt Templates
Prediction Prompt for Llama-8B-Zero models
“Here are all the possible writing intention labels:
Idea Generation: Formulate and record initial thoughts and concepts.
Idea Organization: Select the most useful materials and demarcate those generated ideas in a visually formatted way.
Section Planning: Initially create sections and sub-level structures.
Text Production: Translate their ideas into full languages, either from the writers’ language or borrowed sentences from an external source.
Object Insertion: Insert visual claims of their arguments (e.g., figures, tables, equations, footnotes, itemized lists, etc.).
Cross-reference: Link different sections, figures, tables, or other elements within a document through referencing commands.
Citation Integration: Incorporate bibliographic references into a document and systematically link these references using citation commands.
Macro Insertion: Incorporate predefined commands or packages into a LaTeX document to alter its formatting.
Fluency: Fix grammatical or syntactic errors in the text or LaTeX commands.
Coherence: Logically link (1) any of the two or multiple sentences within the same paragraph; (2) any two subsequent paragraphs; or (3) objects to be consistent as a whole.
Structural: Improve the flow of information by modifying the location of texts and objects.
Clarity: Improve the semantic relationships between texts to be more straightforward and concise.
Linguistic Style: Modify texts with the writer’s writing preferences regarding styles and word choices, etc.
Scientific Accuracy: Update or correct scientific evidence (e.g., numbers, equations) for more accurate claims.
Visual Formatting: Modify the stylistic formatting of texts, objects, and citations.
Identify the most likely next writing intention of a graduate researcher when editing the following LaTex paper draft. Your output should only be a label from the list above.
## Input: before_text
## Output:
Prediction Prompt for Llama-8B-SW models
“Here are all the possible writing intention labels:
Idea Generation: Formulate and record initial thoughts and concepts.
Idea Organization: Select the most useful materials and demarcate those generated ideas in a visually formatted way.
Section Planning: Initially create sections and sub-level structures.
Text Production: Translate their ideas into full languages, either from the writers’ language or borrowed sentences from an external source.
Object Insertion: Insert visual claims of their arguments (e.g., figures, tables, equations, footnotes, itemized lists, etc.).
Cross-reference: Link different sections, figures, tables, or other elements within a document through referencing commands.
Citation Integration: Incorporate bibliographic references into a document and systematically link these references using citation commands.
Macro Insertion: Incorporate predefined commands or packages into a LaTeX document to alter its formatting.
Fluency: Fix grammatical or syntactic errors in the text or LaTeX commands.
Coherence: Logically link (1) any of the two or multiple sentences within the same paragraph; (2) any two subsequent paragraphs; or (3) objects to be consistent as a whole.
Structural: Improve the flow of information by modifying the location of texts and objects.
Clarity: Improve the semantic relationships between texts to be more straightforward and concise.
Linguistic Style: Modify texts with the writer’s writing preferences regarding styles and word choices, etc.
Scientific Accuracy: Update or correct scientific evidence (e.g., numbers, equations) for more accurate claims.
Visual Formatting: Modify the stylistic formatting of texts, objects, and citations.
Identify the most likely next writing intention of a graduate researcher when writing the following LaTex paper draft. Your output should only be a label from the list above.
## Input: before_text
## Output:
Classification Prompt for GPT-4o
“You are a classifier that identify the most likely next writing intention. You will be given a list of all possible writing intention labels with definitions, and an in-progress LaTex paper draft written by a graduate student. Please strictly follow user’s instruction to identify the most likely next writing intention.
Here are the verbalizers of all the possible writing intention labels:
Idea Generation: Formulate and record initial thoughts and concepts.
Idea Organization: Select the most useful materials and demarcate those generated ideas in a visually formatted way.
Section Planning: Initially create sections and sub-level structures.
Text Production: Translate their ideas into full languages, either from the writers’ language or borrowed sentences from an external source.
Object Insertion: Insert visual claims of their arguments (e.g., figures, tables, equations, footnotes, itemized lists, etc.).
Cross-reference: Link different sections, figures, tables, or other elements within a document through referencing commands.
Citation Integration: Incorporate bibliographic references into a document and systematically link these references using citation commands.
Macro Insertion: Incorporate predefined commands or packages into a LaTeX document to alter its formatting.
Fluency: Fix grammatical or syntactic errors in the text or LaTeX commands.
Coherence: Logically link (1) any of the two or multiple sentences within the same paragraph; (2) any two subsequent paragraphs; or (3) objects to be consistent as a whole.
Structural: Improve the flow of information by modifying the location of texts and objects.
Clarity: Improve the semantic relationships between texts to be more straightforward and concise.
Linguistic Style: Modify texts with the writer’s writing preferences regarding styles and word choices, etc.
Scientific Accuracy: Update or correct scientific evidence (e.g., numbers, equations) for more accurate claims.
Visual Formatting: Modify the stylistic formatting of texts, objects, and citations.
Identify the most likely next writing intention of a graduate researcher when editing the following LaTex paper draft. Your output should only be a label from the list above.
## Input: before_text
## Output:
F.1.4 Discussion about Next Intention Prediction Results
According to Table 6, none of them is reaching 0.7 in F1. This is likely due to the intricate nature of the task. The model is asked to predict the next intention by only giving the before text. In the annotation task, the annotator labels the data by looking through multiple consecutive before and after text pairs to determine the current intention rather than looking at the current text to predict the next intention. Another reason is that the next intention chosen by the author does not necessarily mean that it is the only correct intention. We also noticed that BERT and RoBERTa perform much better than the vanilla Llama-8b-Instruct. This is likely due to Llama 8b being under-trained as they have a larger size of parameters, while we fine-tuned it on our data for only one epoch.
F.2 Output Alignment across Writing Sessions
F.2.1 Prompt Templates
Single-intention Session Output Generation
You are an expert writing assistant.
Your role is to improve user-provided text according to specific criteria.
The input will be LaTeX text.
Always output only the improved LaTeX text, with no explanations or notes.
Improve the following LaTeX text according to this criteria:
- {intention}: {intention definition}
Text to improve:
{before_text}
Multi-intention Session Output Generation
You are an expert writing assistant.
Your role is to improve user-provided text according to specific criteria.
The input will be LaTeX text.
Always output only the improved LaTeX text, with no explanations or notes.
Improve the following LaTeX text according to these criteria:
{list of intentions and their definitions}
Text to improve:
{before_text}
F.2.2 Results
Figure 16 shows the relationship between a single-intention session duration and alignment score (either Levenshtein ratio or BERTScore-F1) for two models, GPT-5 and Qwen-2.5-14B.
Figure 17 shows the relationship between a single-intention session duration and alignment score (either Levenshtein ratio or BERTScore-F1) for two models, GPT-5 and Qwen-2.5-14B across different intention categories.
Figure 18 shows the relationship between number of writing intentions in a multi-intention session and alignment score (either Levenshtein ratio or BERTScore-F1) for two models, GPT-5 and Qwen-2.5-14B.
Appendix G More about ScholaWrite Evaluation
We envision ScholaWrite as a valuable resource for training language models and improving future writing assistants for scholarly writing. To evaluate its usability, we conducted experiments training LLMs to mimic the complex, non-linear writing processes of human scholars.
Iteratively generating scholarly writing actions from scratch (called Iterative Self-Writing), mirroring the human writing process: This task focuses on how well the model trained on our dataset can replicate the actual iterative writing and thinking process of scholars, and whether the generated text achieves higher quality compared to LLM-prompted writing.
Setups
During iterative self-writing (Figure 19), a fine-tuned model processes LaTeX-formatted seed document (as “before-text”) with a context prompt to predict the next intention, then revises the text (“after-text”) accordingly given prompt. The revised document then serves as the new seed for the next iteration. This process repeats until a set iteration limit (e.g., 100) is reached. All models use the same train (80%)-test (20%) split across experiments. Figure 23 for training details.
We fine-tune Llama3.1-8B-Instruct (Llama-8B-SW) and compare it to vanilla Llama-8B-Instruct (Llama-8B-Zero) and GPT-4o. Also, seed documents were derived from LaTeX-formatted abstracts of four award-winning NLP papers on diverse topics (Zeng et al., 2024; Lu et al., 2024; Du et al., 2022a; Etxaniz et al., 2024), as shown in Appendix Listings LABEL:table:seed-entry-johnny-LABEL:table:seed-entry-latxa.
Metrics
We evaluated lexical diversity (unique-to-total token ratio), topic consistency (cosine similarity between seed and final output), and intention coverage (unique writing intentions used over 100 iterations). For human evaluation, three native English speakers with LaTeX expertise assessed outputs from Llama-8B-Zero and Llama-8B-SW on accuracy (alignment with predicted intention), alignment (similarity to human writing), fluency(grammatical correctness), coherence(logical structure), and relevance (connection to the seed document) - Refer to Appendix LABEL:sec:appendix:human-eval. Accuracy was judged per iteration, while other metrics used pairwise comparisons. Inter-annotator agreement (IAA)666The IAA scores are 0.84 (SW) and 0.76 (Zero) for accuracy, all 100% for alignment, fluency, coherence, and 49.8% (SW) and 100% (Zero) for relevance. was measured using Krippendorff’s alpha for accuracy and percentage agreement for others.
Llama-8B-SW
Llama-8B-Zero
GPT-4o
| Iter. | Llama-8B-SW | Llama-8B-Zero | GPT-4o |
[…]
Results
Figure 20 shows that Llama-8B-SW consistently produced the most lexically diverse words, generated the most semantically aligned topics (Seeds 1 & 2), and covered the most writing intentions (except Seed 3). These results underscore the value of ScholaWrite in improving scholarly writing quality generated by language models.
However, our human evaluation (Figure 21) revealed that Llama-8B-SW generated less human-like writing, in terms of fluency and logical claims. It also struggled with generating texts aligned with the predicted intentions. See Appendix Tables LABEL:table:human-eval-seed1 to LABEL:table:human-eval-seed4 for more details. Despite the weaknesses, Llama-8B-SW still produced more relevant content (Seed 2), which aligns with topic consistency trends in Figure 20, highlighting the usefulness of ScholaWrite dataset in certain contexts.
Moreover, Llama-8B-SW exhibited the most human-like writing activity patterns over time (Figure 22), which frequently switches between implementation and revision and covers all three high-level processes. Llama-8B-Zero and GPT-4o tend to remain in a single high-level stage throughout all 100 iterations of self-writing (see Appendix LABEL:fig:writing-step-broad-all-model and LABEL:fig:writing-step-intention-all-model for details). Compared to Appendix Figure 13, which depicts frequent transitions across all three stages in an early draft (e.g., the first 100 steps), Llama-8B-SW most closely replicates human writing behaviors in iterative writing tasks. These findings reinforce the potential of ScholaWrite in helping LLMs emulate human scholarly writing processes.
G.1 Training environments
BERT & RoBERTa
We fine-tuned BERT and RoBERTa with the following hyperparameter setups: (1) a learning rate of ; (2) training batch size per device of 8; (3) evaluation batch size per device of 8; (4) the number of training epochs of 10; and (5) a weight decay of . For each model, it took approximately 3.5 hours on one NVIDIA RTX A6000.
Llama
For all experiments, we used baseline models of 4-bit quantized Llama-8B-Instruct777https://huggingface.co/unsloth/Meta-Llama-3.1-8B-Instruct-bnb-4bit, using unsloth library888https://github.com/unslothai/unsloth.
For the ‘after’ text generation subtask from the iterative self-writing experiment, we used the following hyperparameter setups for the fine-tuned Llama-8B-Instruct: (1) only one epoch of training; (2) a weight decay of 0.01; (3) warm-up steps of 10; (4) learning rate of ; and (5) AdamW 8-bit optimizer. Due to computational constraints, we were able to run only one epoch for fine-tuning Llama models on our ScholaWrite dataset. Also, it took approximately 12 hours on one NVIDIA L40s.
GPT-4o
We used the GPT-4o-2024-08-06 version. The iterative writing with 100 iterations took approximately 1 hour on each seed.
G.2 Details About Finetuning Process
The structure of the “after” text differs slightly to help the model learn where and what edits to make. We used the diff_match_patch library to generate a word-level difference array between the “before” and “after” texts. Special tokens (<same>, </same>, <del>, </del>, <add>, </add>) were added to the tokenizer, and the difference array was converted into text wrapped with these tokens. For example, given a “before” text of ‘‘Bad dog’’ and an “after” text of ‘‘Good dog’’, the difference array would be [(-1, ‘Bad’), (1, ‘Good ’), (0, ‘dog’)]. This is converted into: <del>Bad </del><add>Good </add><same>dog</same>. This transformation was applied only to the “after” text, while the “before” text remained as plain LaTeX text.
For finetuning, we randomly split the ScholaWrite dataset into training (80%) and testing (20%) sets. From each intention label in the test set, we sample up to 300 keystroke entries due to budget constraints.
For intention prediction, we fine-tune Llama3.1-8B-Instruct on ScholaWrite training set (Llama-8B-SW-pred”) and compare it to baseline models (Llama-8B-Instruct and GPT-4o) from Section 6.1. For “after-text” generation, we fine-tune another Llama3.1-8B-Instruct model (“Llama-8B-SW-gen”) using the same dataset, with Llama3.1-8B-Instruct and GPT-4o as baselines. The fine-tuning prompt includes task instructions, a verbalizer from human-annotated labels, and “before-text.” While prompts were standardized, task instructions varied by model.
The fine-tuning prompt included task instructions, a verbalizer derived from human-annotated labels, and the “before” text. For fine-tuning, we used QLoRA (Dettmers et al., 2024) to optimize all linear modules of a 4-bit quantized model while maintaining a small memory footprint. Additionally, the embed_tokens and lm_head modules were set as trainable and saved in the final checkpoint. To focus training on the response portion (the “after” text), we used the train_on_response_only function, masking the task instructions, verbalizer, and “before” text with -100s. This ensures the model learns to generate the "after" text without being influenced by instructional input. The model was trained for one epoch with a batch size of 1, 4 gradient accumulation steps, and the AdamW 8-bit optimizer.
During iterative writing, we performed 100 iterations, treating the model’s output under one intention as a single iteration. If the intention predicted by the classification model (fine-tuned Llama3.1-8B-Instruct, as described in Sec 6.1) matched the current predicted intention, the model was prompted to edit the text again. In this case, the newly generated output was not treated as final output in the iteration, and the iteration did not proceed. We moved to the next iteration only when the intention prediction model generated a different intention label than the previous one.
For model setups, we created three pairs of models for each prediction and generation subtasks as follows:
LLama-8B-SW: LLama-8B-Instruct fine-tuned on ScholaWrite dataset (Prediction) & LLama-8B-Instruct fine-tuned on ScholaWrite dataset (Generation), independently
LLama-8B-Instruct: Vanilla LLama-8B-Instruct (Prediction) & Vanilla LLama-8B-Instruct (Generation), independently
GPT-4o: GPT-4o inference (Prediction) & GPT-4o inference (Generation)
Also, due to budget constraints, models had different revision strategies. Llama-8B-SW-* and Llama-8B-Instruct continued revision until the next predicted intention changed. GPT-4o, however, moved to the next iteration regardless. We refer to the fine-tuned Llama-8B model as Llama-8B-ScholaWrite (or Llama-8B-SW) and the vanilla model as Llama-8B-Zero.
G.3 Prompt Templates
G.3.1 Generation Prompt for Llama-8B-Zero models
“You are a computer science researcher with extensive experience of scholarly writing. Here, you are writing a research paper in natural language processing using LaTeX languages.
You currently want to ‘‘put the verbalizer of the predicted intention label’’ (e.g., ‘‘initially create sections and sub-level structures’’ if the predicted label was section planning).
Below is the paper you have written so far. Please strictly follow the writing intention given above and insert, delete, or revise at appropriate places in the paper given below.
Your writing should relate to the paper given below. Do not generate text other than paper content. Do not describe the changes you are making or your reasoning.
## Input: before_text ”
G.3.2 Generation Prompt for Llama-8B-Zero models
“You are a computer science researcher with extensive experience in scholarly writing. Here, you are writing a research paper in natural language processing using LaTeX.
You currently want to ‘‘put the verbalizer of the predicted intention label’’ (e.g., ‘‘initially create sections and sub-level structures’’ if the predicted label was section planning).
Below is the paper you have written so far. Given the paper information below and the corresponding scholarly writing intention, please revise or add to the text to fulfill this writing intention.
You may insert, delete, or revise text at appropriate places in the given paper.
Please provide a complete output. Do not generate text that is nonsensical or unrelated to the given paper information.
## Input: before_text ”
G.3.3 Generation Prompt for GPT-4o
“ You are a computer science researcher with extensive experience of scholarly writing. Here, you are writing a research paper in natural language processing using LaTeX languages.
Your writing intention is to ‘‘put the verbalizer of the predicted intention label’’ (e.g., ‘‘initially create sections and sub-level structures’’ if the predicted label was section planning).
Below is the paper you have written so far. Please strictly follow the writing intention given above and insert, delete, or revise at appropriate places in the paper given below.
Your writing should relate to the paper given below. Do not generate text other than paper content. Do not describe the changes you are making or your reasoning. Do not include sidenotes. Your output should only be the paper draft in latex, without the “‘latex delimiters.
## Input: before_text
## Output:
G.4 Seed Documents for Iterative Self-Writing
We present the four seed documents that we used for the iterative self-writing experiments, as shown in Listings LABEL:table:seed-entry-johnny to LABEL:table:seed-entry-latxa.